### v1

# Not All Clients Are Equal: Personalized Federated Learning on Heterogeneous Multi-Modal Clients

Abstract
Foundation models have shown remarkable capabilities across diverse multi-modal tasks, but their centralized training raises privacy concerns and induces high transmission costs. In contrast, federated learning (FL) offers a distributed alternative without the need to share data. Recently, for the growing demand for personalizing AI models for different user purposes, personalized federated learning (PFL) has emerged. PFL allows each client to leverage the knowledge of other clients for further adaptation to individual user preferences, again without the need to share data. Despite its potential, most PFL studies remain confined to simulated environments, overlooking the data and model heterogeneity that arise in real-world scenarios. In contrast, we first consider large data heterogeneity, evaluating on a new benchmark for multi-modal PFL, spanning 40 distinct tasks with realistic data distribution shifts. We then consider model heterogeneity in that we do not assume that all clients share similar model architectures. To address data heterogeneity, we propose a task-similarity-aware model aggregation method that provides customized global models to each client. For model heterogeneity, we propose a dimension-invariant module that enables knowledge sharing across heterogeneous models. Empirical validations demonstrate that the proposed approach outperforms the state-of-the-art, excelling in both personalization and generalization capabilities.
1 Introduction
Multi-modal Large Language Models (MLLMs) with billions of parameters, such as GPT-4o (hurst2024gpt, ), Qwen2.5-VL (yang2024qwen2, ) and Gemini (team2024gemini, ), have recently demonstrated remarkable performance. This is driven by training on large-scale data collected from various sources using high-performance computing resources, which allows for strong generalizability across diverse multi-modal tasks (chen2024feddat, ). However, such centralized training raises significant challenges, including concerns about data privacy and high transmission costs (alvi2022federated, ; MIR-2022-03-072, ). Moreover, as the demand for personalization grows, further fine-tuning is required to adapt models to individual user preferences (lau2024personalized, ; zhang2024guided, ).
To address both privacy and personalization challenges, personalized federated learning (PFL) (smith2017federated, ) has emerged as a decentralized alternative. The PFL shares locally trained models instead of raw data, while leveraging the knowledge of other clients to further adapt each local model to individual user preferences. Recent PFL methods show that the sharing of knowledge among client models, trained in a distributed manner, significantly improves the performance of personalized MLLMs (scott2024pefll, ; xie2024perada, ).
Figure 1: Overview of Heterogeneous PFL. $L_{i}$ refers to the local model for the $i_{th}$ client. For each client, new data are encountered continuously in a streaming manner, with each client focusing on a different task (i.e., data heterogeneity). In addition to data heterogeneity, model architectures may differ across clients (i.e., model heterogeneity) due to differences in hardware constraints.
Despite its potential, most PFL studies remain confined to simulated environments, as they often overlook the client heterogeneity present in real-world scenarios (bujotzek2025real, ; zhang2024fedcvd, ). Specifically, in real-world scenarios, clients typically have distinct models tailored to their environments (i.e., model heterogeneity) and train on highly personalized, client-specific data (i.e., data heterogeneity). Although recent work has begun to address client heterogeneity, it mostly considers either model heterogeneity (fang2023robust, ; wu2024fiarse, ; yi2023fedgh, ) or data heterogeneity (chen2024feddat, ; xie2024perada, ; Tamirisa_2024_CVPR, ). Some attempt to tackle both, but often in simplified setups, e.g., assigning each client a model with different ranks of LoRA (hu2022lora, ) adapter while keeping the base architecture identical (cho2024heterogeneous, ; bai2024federated, ), or simulating data heterogeneity by splitting a single dataset into non-i.i.d. partitions and assigning them to different clients (li2019fedmd, ; alam2022fedrolex, ).
To bridge the gap between simulated and real-world federated learning, we propose FedMosaic, a novel multi-modal PFL method that jointly addresses data and model heterogeneity. We reflect real-world data heterogeneity by assigning each client a distinct multi-modal task, e.g., visual question answering and visual reasoning, rather than focusing different clients on non-i.i.d. splits of a single image classification dataset. Moreover, we also account for distribution shifts inherent in the real world, where each client continuously encounters new data over time. To support this, we propose DRAKE, a comprehensive benchmark for multi-modal PFL with 40 diverse tasks. To the best of our knowledge, this is the first benchmark for multi-modal FL that considers data heterogeneity, as well as distribution shifts. We provide an overview of the heterogeneous PFL scenario in Fig. 1.
We then explore effective knowledge sharing across clients that addresses heterogeneous tasks. Previous work shows that naive model averaging (mcmahan2017communication, ) can degrade performance (wu2024personalized, ) due to interference between models trained on heterogeneous tasks (yadav2023ties, ). To reduce interference, inspired by models trained on similar tasks having less conflict (padmakumar2022exploring, ; gurulingan2022curbing, ), we propose SIMilarity-aware Aggregation (SIMA). SIMA constructs a customized global model for each client, based on task similarity. This enables related clients to share knowledge more effectively.
Aggregating model weights enables knowledge sharing without sharing raw data, but only works in the case of identical model architectures between clients (fan2024fedtsa, ). In practice, clients often use different models based on individual hardware constraints or task complexities. Although federated distillation (FD) approaches have been proposed to aggregate heterogeneous models’ knowledge using logits from public data (xie2024perada, ; li2024federated, ), domain discrepancies between public and client data hinder effective knowledge transfer (zhang2023towards, ; wang2023towards, ). Moreover, logit extraction incurs substantial computational costs (malladi2023fine, ), especially for large models. Instead, we introduce PQ-LoRA, which incorporates dimension-invariant modules $P\in\mathbb{R}^{r\times r}$ and $Q\in\mathbb{R}^{r}$, whose dimensions depend only on low-rank size $r$, while independent of hidden dimension size, making it shareable among heterogeneous models.
We summarize our contributions as follows:
  • Proposing DRAKE, a comprehensive multi-modal federated learning benchmark.
  • Proposing SIMA, a model aggregation strategy that promotes selective knowledge sharing among similar-task models only, addressing data heterogeneity.
  • Proposing PQ-LoRA, shareable across heterogeneous models, addressing model heterogeneity.
2 Related Work
Personalized Federated Learning.
Federated learning aims to train a strong global model in a distributed manner while preserving privacy by sharing model weights instead of raw data (yurdem2024federated, ; hu2024overview, ). However, global models often fail to converge with heterogeneous client data distributions (i.e., non-i.i.d.) (li2020federated, ; tamirisa2024fedselect, ). To address this challenge, personalized federated learning (PFL) (smith2017federated, ) has emerged alongside the growing importance of model personalization, which allows large foundation models to adapt to individual user preferences, contexts, and needs (shashidhar2024unsupervised, ; zhang2024personalization, ). PFL aims to train a personalized model on each client’s local data by leveraging shared knowledge from other clients to enhance both personalization and generalization (xie2024perada, ).
Data Heterogeneity.
Prior work simulates data heterogeneity by partitioning a single image classification dataset (e.g., MNIST (deng2012mnist, )) into multiple non-i.i.d. subsets, assigning each to a different client. However, assuming such label skew fails to capture real-world data heterogeneity (borazjani2025redefining, ), where clients often tackle different tasks across diverse vision and language domains (madni2024exploiting, ). While  (chen2024feddat, ) makes progress by assigning different VQA datasets to clients, that work remains limited to a single task type (i.e., VQA) and single-image inputs. In contrast, our proposed benchmark, DRAKE, incorporates a broader range of multi-modal tasks, including VQA, visual reasoning, and visual relation detection, covering single- and multi-image inputs, while also modeling temporal distribution shifts within each client, reflecting the evolving and non-stationary nature of real-world data (he2020incremental, ; garg2024tic, ).
Model Heterogeneity.
Federated distillation approaches aggregate knowledge of heterogeneous models using logits from public data (xie2024perada, ; fan2025fedmkt, ). FedMD (li2019fedmd, ) and PerAda (xie2024perada, ) distill averaged logits from local models, while FedMKT (fan2025fedmkt, ) uses those with the lowest loss. However, domain gaps between public and client data hinder knowledge transfer (zhang2023towards, ; wang2023towards, ), while logit sharing raises privacy risks (sun2021vertical, ; lyu2022privacy, ) and incurs significant computational overhead for large models (malladi2023fine, ).
To address the limitation of federated distillation, recent works propose direct model aggregation for heterogeneous models under LoRA-based fine-tuning. HETLoRA (cho2024heterogeneous, ) aligns LoRA modules with varying ranks through zero-padding during aggregation, followed by truncation for redistribution. Similarly, FlexLoRA (bai2024federated, ) multiplies client-specific LoRA matrices to form full matrices, aggregates them, and applies SVD to distribute the aggregated matrix according to each client’s local rank. They address the varying rank sizes, but assume (i) the same input/output dimensions and (ii) identical depths across client models. However, heterogeneous architectures typically have different hidden dimensions and depths (yoshino2023spatially, ; yao2024exploring, ). In contrast, our proposed PQ-LoRA accommodates both different hidden dimensions and varying depths, making it suitable for a more general heterogeneous setup.
3 Preliminaries
3.1 Low-Rank Adaptation (LoRA)
LoRA (hu2022lora, ) is a widely used parameter-efficient tuning method. It assumes that fine-tuning updates lie in a low-rank space. Building on this assumption, LoRA constrains the weight update $\Delta W$ for a pre-trained weight matrix $W_{p}\in\mathbb{R}^{d_{O}\times d_{I}}$ through low rank decomposition using matrices $A\in\mathbb{R}^{r\times d_{I}}$ and $B\in\mathbb{R}^{d_{O}\times r}$, where the rank $r\ll\min(d_{O},d_{I})$. During training, only $A$ and $B$ are updated, while $W_{p}$ remains frozen. As a result, when fine-tuning with LoRA, the original output $h_{O}=W_{p}h_{I}\in\mathbb{R}^{d_{O}}$ for an input hidden state $h_{I}\in\mathbb{R}^{d_{I}}$ is modified by incorporating the low-rank update as:
$h_{O}=(W_{p}+\Delta W)h_{I}=(W_{p}+BA)h_{I}.$ (1)
3.2 Problem Statement of Personalized Federated Learning
We consider a PFL setup with $N$ clients, where each client $i\in[N]$ has a local dataset $\mathcal{D}_{i}$. Reflecting real-world scenarios where data arrives incrementally (seo2024learning, ; seo2025budgeted, ), we assume that each client receives a continuous stream of samples $(x^{(i)}_{1},y^{(i)}_{1}),(x^{(i)}_{2},y^{(i)}_{2}),\cdots$. For client models, given a set of model architectures $W=\{W_{1},\dots,W_{K}\}$, client $i$ selects its local pre-trained model $W_{(p,i)}=V(i)$, based on its hardware constraints, where $V$ is a mapping function $V:\{1,\dots,N\}\rightarrow W$. To improve training and communication efficiency, each client trains and shares only its local LoRA adapter $L_{i}$, while freezing the pre-trained model $W_{(p,i)}$, as described in Sec. 3.1. Let $f(W_{(p,i)},L_{i},x)$ denote the forward pass that outputs the logit vector for input $x$, and let the loss function be $\ell(f(W_{(p,i)},L_{i},x),y)$. The empirical loss for client $i$ over its local dataset $\mathcal{D}_{i}$ can be defined as $\mathcal{J}_{i}(L_{i})=\frac{1}{|\mathcal{D}_{i}|}\sum_{(x,y)\in\mathcal{D}_{i}}\ell(f(W_{(p,i)},L_{i},x),y).$ The objective of PFL is then formulated as:
$\min_{\{L_{1},\dots,L_{N}\}}\frac{1}{N}\sum_{i=1}^{N}\mathcal{J}_{i}(L_{i}).$ (2)
4 Proposed Method
Reflecting the real-world heterogeneity in PFL, we consider both data heterogeneity and model heterogeneity. To address both heterogeneity, we propose FedMosaic, which consists of two key components: SIMA (SIMilarity-aware Aggregation) and PQ-LoRA. SIMA addresses data heterogeneity by reducing interference between local models trained on heterogeneous tasks during aggregation via knowledge sharing among similar-task models only. However, model aggregation (i.e., merging by parameter averaging) becomes infeasible when clients have different architectures. To address model heterogeneity, we introduce PQ-LoRA, which incorporates shareable modules $P\in\mathbb{R}^{r\times r},Q\in\mathbb{R}^{r}$, whose dimensions depend only on low-rank size $r$, while independent of hidden dimension size. We provide an overview of FedMosaic in Fig. 2 and a pseudocode in Sec. A.21.
Figure 2: Overview of our proposed FedMosaic. $L_{i}$ and $G_{i}$ refer to the local and the customized global adapter for the $i_{\text{th}}$ client, respectively. $\hat{g}_{i}$ denotes the decayed last-layer gradient for the $i_{\text{th}}$ client, and $|W|$ the number of layers in model $W$. $\beta_{l}$ is a learnable gating parameter that balances the output from the global and local models in the $l$-th layer. $h$ and $W_{p}$ denote the hidden state input and the pre-trained weight, respectively. $G$ is updated every round and kept frozen during local training.
4.1 Similarity-Aware Aggregation
Model aggregation enables building a single model that excels across multiple tasks without accessing raw data (yadav2024matters, ; wei2025modeling, ). It is thus widely used to construct a shared global model in federated learning. However, aggregating models trained on different tasks often causes parameter interference (yadav2023ties, ; yang2024model, ). Recent work shows models solving similar tasks share more transferable knowledge with fewer conflicts (gurulingan2022curbing, ; padmakumar2022exploring, ). Motivated by this, rather than averaging local models into a single global model, we aggregate them based on task similarity, assigning higher weights to models solving similar tasks and lower weights to less related ones, providing each client with a customized global model.
Client-Wise Gradient $g_{i}$.
To measure task similarity between clients, we calculate the similarity of client-wise gradients. Formally, we calculate the gradient for $i_{\text{th}}$ client as follows:
$g_{i}=\mathbb{E}_{z\subset\mathcal{D}_{i}}\left[\nabla_{W_{s}}\ell(z)\right],$ (3)
where $\mathcal{D}_{i}$ is the data stream of $i_{\text{th}}$ client seen up to this point, $\ell$ refers to the loss function, $z\subset\mathcal{D}_{i}$ is the mini-batch in the data stream, and $W_{s}$ is a small-scale frozen pre-trained model $W_{s}$. For computational efficiency, we: (1) use gradients from a small-scale frozen pre-trained model $W_{s}$ that provides sufficient representativeness with reduced overhead (lee2024concept, ; maharana2025adaptinfty, ), and (2) compute only the last-layer gradient, as preceding layers’ gradients are proportional to it based on the chain rule (koh2023online, ; seo2025budgeted, ). Moreover, we compute the gradient $g_{i}$ in every $m$ iterations rather than in every batch, thus incurring negligible additional cost, as shown in Sec. A.6. Note that we measure gradients of a frozen pre-trained model, not the training model. This is because in heterogeneous PFL, clients train on diverse tasks, such as VQA and visual reasoning, and gradient similarities from models trained on heterogeneous data may not capture actual similarity (tang2020similarity, ; evans2024data, ).
Decayed Client-Wise Gradient $\hat{g}_{i}$.
However, $g_{i}$ may fail to reflect task similarity under shifting data distributions, as it is an expectation over the entire data stream $D_{i}$ (Eq. 3), i.e., unweighted average of gradients across all time points, ignoring forgetting of the model over time. Consequently, client 1 learning $A\rightarrow B$ and client 2 learning $B\rightarrow A$ yield the same $g_{i}$, despite retaining different knowledge due to catastrophic forgetting (McCloskey1989catastrophic, ; ratcliff1990connectionist, ). To reflect the model knowledge shifts under distribution shifts, we introduce decayed gradient similarity $\hat{g}_{i}$, computed using exponential moving average (EMA) of past gradients, inspired by the exponential decay of knowledge in forgetting (mahto2021multi, ; chien2021slower, ; seo2025budgeted, ). Formally, for the $i_{\text{th}}$ client, $\hat{g}_{i}(t)$ with EMA ratio $\alpha$ at timestep $t$ is defined recursively as:
$\hat{g}_{i}(t)=(1-\alpha)*\hat{g_{i}}(t-1)+\alpha*g_{i}(t),$ (4)
where $g_{i}(t)=\nabla_{W_{s}}\ell(z_{t})$ is the gradient vector for the given batch $z_{t}\subset\mathcal{D}_{i}$ at timestep $t$.
Using the decayed client-wise gradients $\{\hat{g}_{1},\dots,\hat{g}_{N}\}$ from $N$ clients, we construct a client-similarity matrix $S\in\mathbb{R}^{N\times N}$, where $S_{ij}=\text{cos}(\hat{g}_{i},\hat{g}_{j})$. The customized global module for the $i_{\text{th}}$ client, $G_{i}$, is then constructed by weighted aggregation of local modules =$\mathcal{L}=\{L_{1},\dots,L_{N}\}$ as follows:
$G_{i}=\sum_{j=1}^{N}w_{ij}~{}L_{j},~{}~{}~{}~{}w_{ij}=\frac{e^{{\text{cos}(\hat{g}_{i},\hat{g}_{j})}/\tau}}{\sum_{n=1}^{N}e^{\text{cos}(\hat{g}_{i},\hat{g}_{n})/\tau}},$ (5)
where $\tau$ denotes the softmax temperature and cos denotes the cosine similarity.
4.2 PQ-LoRA
LoRA matrices $A\in\mathbb{R}^{r\times d_{I}}$ and $B\in\mathbb{R}^{d_{O}\times r}$ depend on model-specific hidden dimensions $d_{I}$ and $d_{O}$, and thus cannot be directly aggregated across different models. To share knowledge among heterogeneous architectures, we introduce PQ-LoRA, which inserts dimension-invariant modules $P\in\mathbb{R}^{r\times r}$ and $Q\in\mathbb{R}^{r}$ between $A$ and $B$. The dimensions of these modules depend only on the low-rank size $r$, making them shareable across heterogeneous architectures. We illustrate a comparison of PQ-LoRA with conventional LoRA in Fig. 3. Formally, PQ-LoRA outputs $h_{O}$ as:
$h_{O}=W_{p}h_{I}+B(PAh_{I}+Q),$ (6)
for input hidden states $h_{I}$ and pre-trained weight $W_{p}$. After local training on each client, the $P$ and $Q$ modules are aggregated and shared between heterogeneous clients for knowledge sharing.
Figure 3: The illustration of (a) Conventional LoRA and (b) PQ-LoRA. While $A$ and $B$ are trainable in conventional LoRA, PQ-LoRA freezes both, updating only the dimension-invariant modules $P\in\mathbb{R}^{r\times r}$ and $Q\in\mathbb{R}^{r}$ during training.
Although $P$ and $Q$ are shareable across heterogeneous architectures having different dimensions, two main challenges hinder their aggregation: (i) depth heterogeneity - heterogeneous architectures often differ in depth, making it non-trivial to determine which layers across models should be aggregated; and (ii) weight misalignment - interpolating weights with different optimization trajectories can degrade performance (entezari2021role, ; jordanr2023epair, ; neyshabur2020being, ; stoica2025model, ). To address these challenges, we adopt two strategies: (i) block-wise aggregation, which aligns layers at the same relative depth, and (ii) weight alignment, ensuring heterogeneous models share the same initialization. For simplicity, we consider aggregating two heterogeneous models, though it naturally extends to multiple heterogeneous models.
4.2.1 Block-Wise Aggregation
We measure layer-wise representation alignment between two MLLMs, $W_{i}$ and $W_{j}$, with different depths using CKA (kornblith2019similarity, ), and observe high similarity between layers at similar relative depths, i.e., approximately linear alignment (Sec. A.7). Accordingly, we divide each model into $N_{B}$ blocks linearly and attach a PQ-LoRA to each block’s final layer, enabling cross-model sharing at relevant depths. $I_{k}$, the attachment layer index of the $k$-th PQ-LoRA in a $|W|$-layer model, is defined as:
$I_{k}=\begin{cases}k\cdot\left\lfloor\frac{|W|}{N_{B}}\right\rfloor&\text{for}\quad k=1,2,\ldots,N_{B}-1\\ |W|&\text{for}\quad k=N_{B}.\end{cases}$ (7)
4.2.2 Weight Alignment of PQ-LoRA
We then align the $N_{B}$ number of PQ-LoRA modules at layer indices $\mathcal{I}_{i}=\{I^{i}_{1},\dots,I^{i}_{N_{B}}\}$ in model $M_{i}$ and $\mathcal{I}_{j}=\{I^{j}_{1},\dots,I^{j}_{N_{B}}\}$ in model $M_{j}$. For simplicity, we describe the alignment for a single pair $X=(\{P_{i},Q_{i},A_{i},B_{i}\},\{P_{j},Q_{j},A_{j},B_{j}\})$, which is generalized across all $N_{B}$ pairs.
Motivated by findings that models fine-tuned from the same initialization share optimization paths and can be merged without interference (wortsman2022model, ; wortsman2022robust, ; yadav2023ties, ), we set the pair $X$ to share a common initialization. Dimension-invariant modules, $P_{i},P_{j}\in\mathbb{R}^{r\times r}$ and $Q_{i},Q_{j}\in\mathbb{R}^{r}$, can share the same weight directly, while dimension-dependent modules, $A_{i}\in\mathbb{R}^{r\times d_{i}}$, $A_{j}\in\mathbb{R}^{r\times d_{j}}$ and $B_{i}\in\mathbb{R}^{d_{i}^{\prime}\times r}$, $B_{j}\in\mathbb{R}^{d_{j}^{\prime}\times r}$, cannot due to dimension mismatches. Although $P$ and $Q$ are only shared between heterogeneous models $M_{i}$ and $M_{j}$, $A$ and $B$ should also be aligned, as they affect the output $h_{O}$ (Eq. 6), thereby affecting optimization trajectories. Therefore, we align $A_{i},A_{j}$ and $B_{i},B_{j}$, ensuring aligned representations for the same input, even in parameter spaces with different dimensions.
Aligning $A$ Matrices.
We first align $A_{i}\in\mathbb{R}^{r\times d_{i}}$ and $A_{j}\in\mathbb{R}^{r\times d_{j}}$ using MSE loss in the shared $r$-dimensional space with publicly available data $\mathcal{D}_{p}$. For $\mathcal{D}_{p}$, we use a subset of pre-training data of the employed MLLMs, rather than any downstream client data. In particular, we use the smaller model $A_{i}$ (where $d_{i}<d_{j}$) as the pivot model by freezing it and updating the larger model $A_{j}$ to minimize the L2 loss on $r$-dimensional representations obtained from public data $\mathcal{D}_{p}$, as:
$\min_{A_{j}}\frac{1}{|\mathcal{D}_{p}|}\sum_{(x,y)\in\mathcal{D}_{p}}\text{MSE}(A_{i}(x),A_{j}(x)),$ (8)
where $A(x)$ denotes $r$-dimensional feature extracted by $A$ for input $x$.
Aligning $B$ Matrices.
We then align $B_{i}$ and $B_{j}$. Since they have different output dimensions $d^{\prime}_{i}\neq d^{\prime}_{j}$, MSE loss is inapplicable. Instead, we employ canonical correlation analysis (CCA) (DeBie2005, ), which finds projection matrices that maximize the correlation between two feature sets. Using $m$ public data $\{(x_{k},y_{k})\}_{k=1}^{m}\subset D_{p}$, we extract output features $H_{i}\in\mathbb{R}^{m\times d^{\prime}_{i}}$ and $H_{j}\in\mathbb{R}^{m\times d^{\prime}_{j}}$ from $B_{1}$ and $B_{2}$, then find projection matrices $\Pi_{i}$ and $\Pi_{j}$, which project them to the maximally correlated space. We project $B_{i}$ to the shared maximally correlated space, i.e., $(\Pi_{i})^{T}\cdot B_{i}$, and subsequently bring it to $B_{j}$ space by inverse projection, as $B_{j}=({\Pi_{j}}^{-1})^{T}\cdot(\Pi_{i})^{T}\cdot B_{i}$.
Further Details.
After aligning PQ-LoRA, we update shareable modules $P$ and $Q$, while freezing $A$ and $B$ during local training to preserve alignment. We provide theoretical justification for this freezing approach in Sec. A.1. Moreover, during alignment, we enforce orthogonality in $A$ and $B$ to maximize the expressive capacity (i.e., span) of PQ-LoRA weight updates, following Theorem 1. We provide details of enforcing orthogonality in Sec. A.4 and the proof of Theorem 1 in Sec. A.2.
Theorem 1.
If the column vectors of matrix $B\in\mathbb{R}^{d_{O}\times r}$ are orthogonal and the row vectors of matrix $A\in\mathbb{R}^{r\times d_{I}}$ are orthogonal, then the span of the weight update space of PQ-LoRA, $\text{span}\{\Delta W\}$, has $r^{2}$ dimension, which is the maximum possible dimension under frozen $B$ and $A$.
After receiving the aggregated PQ-LoRA (i.e., global PQ-LoRA) by SIMA at each communication round, clients freeze it during training to preserve global knowledge and update only the local model for personalization. Specifically, at the $l_{\text{th}}$ layer, given an input hidden state $h_{I}$, we combine the output of the local LoRA (i.e., $h_{L}$), the frozen global LoRA (i.e.,$h_{G}$), and the pre-trained weights (i.e., $W_{p}h_{I}$), by adaptively balancing them using a learnable gating parameter $\beta$. With sigmoid-normalized balancing parameter $\tilde{\beta}=\sigma(\beta)$, the output hidden state $h_{O}$ is computed as follows:
$h_{O}=W_{p}h_{I}+(1-\tilde{\beta})h_{L}+\tilde{\beta}h_{G}.$ (9)
5 DRAKE Benchmark
We propose a comprehensive multi-modal FL benchmark, called DRAKE, with three key merits: (i) Task heterogeneity: Each client handles distinct multi-modal tasks (e.g., visual reasoning, VQA), while existing benchmarks merely assign non-i.i.d. subsets split from a single dataset. (ii) Dynamic distribution: Client datasets contain progressive tasks (e.g., encountering new visual concepts), simulating real-world distribution shifts. To our knowledge, DRAKE is the first benchmark that supports the multi-modal federated continual learning scenario. (iii) Evaluation for generalizability: DRAKE incorporates unseen task data to evaluate the generalization and fast adaptation capabilities of FL models. We summarize the comparison with existing FL benchmarks in Tab. 1.
As shown in Fig. 4, DRAKE consists of three training task subgroups, i.e., VQA, visual relation, and multi-modal reasoning, and two unseen task subgroups, i.e., new task types and new visual domains. Specifically, it comprises 40 heterogeneous tasks sourced from 19 different multi-modal datasets, totaling 375k images and 274k questions. We provide more details of DRAKE in Sec. A.18.
Dataset Multi-Image Support Multi-Data Sources Distribution Shifts Multi- Modalities Unseen Evaluation
Split-CIFAR (ICML 2021)
NonIID-50 (ICML 2021)
Leaf-FCL (ICLR 2023)
MNIST-Shuffle (ICLR 2024)
HC-FMTL (CVPR 2024)
Profit (NeurIPSW 2023) -
FEDLEGAL (ACL 2023) -
Fed-Aya (NeurIPS 2024) -
FedDPA (NeurIPS 2024) -
HFLB (AAAI 2024)
DRAKE (Ours)
Table 1: Comparison of FL benchmarks across key dimensions: Multi-Image Support (handling multiple images per input), Multi-Data Sources (using diverse datasets vs. non-i.i.d. splits of single dataset), Distribution Shifts (evolving client data distributions), Multi-Modalities (single or multiple data types), and Unseen Evaluation (testing on tasks unseen during training).
Figure 4: Overview of DRAKE.
6 Experiments
6.1 Setups
Models.
To simulate model heterogeneity in federated MLLM training, we employ LLaVA-1.5 (liu2023visual, ) variants. Specifically, for LLM of LLaVA, we employ various sizes of Llama-3 (grattafiori2024llama, ) (i.e., Llama-3.2-1B, Llama-3.2-3B, Llama-3.1-8B) and Qwen-2.5 (yang2024qwen2, ) (i.e., Qwen2.5-0.5B, Qwen2.5-1.5B, Qwen2.5-3B). For image encoder, we use CLIP ViT-L/336px (radford2021learning, ) in all variants, following  (radford2021learning, ; zhang2025llava, ).
Baselines.
We compare FedMosaic with state-of-the-art PFL methods. DITTO (li2021ditto, ), FedSim (pillutla2022federated, ), FedIT (zhang2024towards, ), TAKFL (morafah2024towards, ), FedDPA (long2024dual, ), FedDAT (chen2024feddat, ), PerAda (xie2024perada, ), and FedMKT (fan2025fedmkt, ). We also compare with supervised fine-tuning (SFT), which trains each client model independently without knowledge sharing and often performs better than participating in FL under data heterogeneity (ghari2024personalized, ).
Federated distillation baselines (i.e., TAKFL, PerAda, and FedMKT) share knowledge between heterogeneous models via logits, whereas FedMosaic employs PQ-LoRA. For the other baselines, we combine them with Fed-ET (cho2022heterogeneous, ), which aggregates models among homogeneous clients.
Metrics.
We report $A_{last}$, the accuracy at the end of training, and $A_{\text{AUC}}$ (koh2022online, ), which computes the area under the accuracy curve by measuring accuracy at each evaluation period to capture intermediate performance. For all experiments, we set the evaluation period to every five rounds and run them with 3 different random seeds, reporting the average performance along with the standard deviation.
Benchmarks.
We evaluate FedMosaic on our proposed DRAKE and HFLB (chen2024feddat, ), a multi-modal PFL benchmark. Our evaluation covers PFL-Static (i.e., i.i.d. client data distributions) and the more realistic PFL-Dynamic setup (i.e., intra-client distribution shifts with four incremental tasks). To evaluate the PFL-Dynamic setup on HFLB dataset, we partition the client data by question types (e.g., color, shape, count, and size). In addition, we evaluate on LLM-FL benchmarks, i.e., Fed-LLM (ye2024fedllm, ) and Fed-Aya (singh2024aya, ). See Sec. A.5 for details of the benchmarks and implementation details.
6.2 Quantitative Analysis
In all experiments, we evaluate each client’s model on its own task (i.e., ‘Self’) and on tasks from other clients (i.e., ‘Others’), reporting the average performance across clients. ‘Self’ performance shows personalization, while ‘Others’ indicates generalizability. Although personalization is the main goal of PFL, generalization ability is also crucial for continual personalization, as it enables rapid adaptation to newly arriving tasks in the future and addition of new clients during training (finn2017model, ; RAO2023271, ).
Method DRAKE-Dynamic HFLB-Dynamic
Self Others Self Others
$A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$
SFT 65.79$\pm$0.20 57.62$\pm$0.26 47.66$\pm$0.13 46.67$\pm$0.11 80.65$\pm$0.41 77.79$\pm$0.08 61.49$\pm$0.09 61.17$\pm$0.13
DITTO (ICML 2021) 59.91$\pm$0.18 54.16$\pm$0.06 47.45$\pm$0.36 46.70$\pm$0.10 78.92$\pm$0.12 76.33$\pm$0.13 61.42$\pm$0.59 59.57$\pm$1.84
FedSim (ICML 2022) 63.98$\pm$1.19 56.42$\pm$0.70 47.04$\pm$0.16 46.15$\pm$0.17 79.30$\pm$0.26 76.60$\pm$0.25 59.48$\pm$0.80 58.84$\pm$0.81
FedIT (ICASSP 2024) 66.11$\pm$0.27 57.86$\pm$0.27 47.63$\pm$0.16 46.62$\pm$0.04 79.90$\pm$0.38 77.21$\pm$0.50 60.29$\pm$1.47 60.37$\pm$1.25
TAKFL (NeurIPS 2024) 64.54$\pm$0.85 56.19$\pm$0.98 47.38$\pm$0.06 46.47$\pm$0.14 79.23$\pm$0.35 76.75$\pm$0.46 61.04$\pm$0.62 59.62$\pm$1.91
FedDPA (NeurIPS 2024) 63.34$\pm$0.23 55.74$\pm$0.38 47.61$\pm$0.21 46.64$\pm$0.20 80.42$\pm$0.41 77.67$\pm$0.04 61.70$\pm$0.08 61.10$\pm$0.09
FedDAT (AAAI 2024) 58.47$\pm$1.10 54.38$\pm$0.28 48.91$\pm$0.42 47.78$\pm$0.18 78.98$\pm$0.38 76.45$\pm$0.62 64.82$\pm$0.54 64.16$\pm$0.39
PerAda (CVPR 2024) 59.75$\pm$1.06 54.48$\pm$0.13 47.30$\pm$0.45 46.72$\pm$0.03 78.61$\pm$0.66 75.92$\pm$0.42 61.24$\pm$0.95 60.83$\pm$0.06
FedMKT (COLING 2025) 61.38$\pm$0.18 55.45$\pm$0.31 47.50$\pm$0.07 46.68$\pm$0.08 78.70$\pm$0.38 75.94$\pm$0.68 61.12$\pm$0.26 59.59$\pm$1.86
FedMosaic (Ours) 67.86$\pm$0.51 59.83$\pm$0.16 51.16$\pm$0.04 49.36$\pm$0.08 80.90$\pm$0.30 79.18$\pm$1.04 66.66$\pm$0.77 65.80$\pm$0.30
Method DRAKE-Static HFLB-Static
Self Others Self Others
$A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$
SFT 68.50$\pm$0.42 63.84$\pm$0.27 47.91$\pm$0.39 47.56$\pm$0.23 80.33$\pm$1.02 79.46$\pm$0.27 61.38$\pm$0.96 61.19$\pm$0.37
DITTO (ICML 2021) 63.67$\pm$1.50 59.17$\pm$1.25 48.22$\pm$0.09 47.65$\pm$0.03 79.65$\pm$0.87 78.23$\pm$0.25 61.98$\pm$0.37 61.38$\pm$0.08
FedSim (ICML 2022) 66.75$\pm$0.56 61.70$\pm$0.60 46.93$\pm$0.25 46.76$\pm$0.02 79.17$\pm$0.16 78.48$\pm$0.53 59.22$\pm$0.59 59.45$\pm$0.44
FedIT (ICASSP 2024) 68.71$\pm$0.04 63.89$\pm$0.55 47.91$\pm$0.22 47.60$\pm$0.17 80.08$\pm$1.07 79.41$\pm$0.17 61.32$\pm$0.75 61.16$\pm$0.27
TAKFL (NeurIPS 2024) 67.28$\pm$0.01 62.42$\pm$0.36 47.55$\pm$0.23 47.51$\pm$0.07 79.45$\pm$0.33 78.42$\pm$0.99 60.61$\pm$1.63 60.45$\pm$1.31
FedDPA (NeurIPS 2024) 66.09$\pm$1.51 61.36$\pm$0.14 47.93$\pm$0.30 47.69$\pm$0.11 79.60$\pm$1.37 78.55$\pm$0.50 61.33$\pm$0.25 60.94$\pm$0.01
FedDAT (AAAI 2024) 61.28$\pm$0.07 57.92$\pm$0.02 49.37$\pm$0.02 48.78$\pm$0.04 79.17$\pm$0.37 77.72$\pm$0.71 65.86$\pm$0.37 64.79$\pm$0.23
PerAda (CVPR 2024) 63.67$\pm$1.01 59.03$\pm$1.19 48.13$\pm$0.03 47.60$\pm$0.05 79.64$\pm$0.70 78.24$\pm$0.17 61.93$\pm$0.38 61.36$\pm$0.07
FedMKT (COLING 2025) 65.81$\pm$1.09 60.37$\pm$1.16 47.19$\pm$0.86 46.44$\pm$1.49 80.64$\pm$0.50 78.48$\pm$0.56 61.15$\pm$0.50 61.11$\pm$0.28
FedMosaic (Ours) 70.10$\pm$0.53 64.24$\pm$0.40 51.57$\pm$0.24 50.42$\pm$0.12 81.63$\pm$0.24 80.11$\pm$0.12 67.03$\pm$0.01 66.02$\pm$0.29
Table 2: Quantitative comparison in heterogeneous PFL. ‘Self’ denotes evaluation on a client’s own data, while ‘Others’ denotes evaluation on data from other clients. In DRAKE, 4 clients use LLaVA-Llama3.2-1B models and 6 clients use LLaVA-Llama3.2-3B models, while in HFLB, 3 clients use LLaVA-Llama3.2-1B models and 6 clients use LLaVA-Llama3.2-3B models. SFT refers to supervised fine-tuning on each client’s data without cross-client knowledge sharing.
Method $A_{last}$ / $A_{\text{AUC}}$
LLaVA-Llama3.2-1B LLaVA-Llama3.2-3B
Client 1 Client 2 Client 3 Client 4 Client 5 Client 6 Client 7 Client 8 Client 9 Client 10
SFT 76.45 / 59.87 58.10 / 54.35 66.63 / 59.57 62.85 / 53.16 62.00 / 61.51 76.56 / 66.88 62.59 / 54.63 68.54 / 59.65 65.92 / 51.93 58.30 / 54.61
FedMosaic (Ours) 77.42 / 60.45 58.50 / 55.26 66.71 / 59.60 63.13 / 55.84 68.87 / 63.60 77.61 / 69.80 69.19 / 56.66 70.54 / 58.93 70.68 / 54.09 59.21 / 55.32
Difference +0.97 / +0.57 +0.40 / +0.91 +0.08 / +0.04 +0.27 / +2.67 +6.86 / +2.09 +1.05 / +2.93 +6.61 / +2.03 +2.00 / -0.72 +4.76 / +2.16 +0.91 / +0.71
Table 3: Per-client performance of FedMosaic vs. SFT in a heterogeneous PFL-Dynamic setup on DRAKE. The Difference row shows the performance gain/loss of FedMosaic compared to SFT. Clients 1-4 use LLaVA-Llama3.2-1B, while Clients 5-10 use LLaVA-Llama3.2-3B.
Heterogeneous Multi-Modal Clients.
We first evaluate FedMosaic in multi-modal heterogeneous PFL-Dynamic and -Static setups, where clients employ different model architectures (LLaVA-Llama3-1B or 3B). As shown in Tab. 2, FedMosaic outperforms baselines on both clients’ own tasks (‘Self’) and others’ tasks (‘Others’) under both static and dynamic distributions. HFLB’s single VQA task with single-image inputs converges faster than DRAKE, leading to smaller gaps among baselines.
Note that the results in Tab. 2 report the average performance across all clients. Prior work has shown that local client training (i.e., SFT) may be sufficient for personalization in some tasks (wozniak2024personalized, ; mosbach2021on, ), which can make the overall average improvements appear small. For instance, as shown in the per-client accuracy in Tab. 3, FedMosaic shows only marginal improvements over SFT for clients 2 and 3. In contrast, for clients 5, 7, and 9, FedMosaic significantly improves personalization by sharing knowledge. We provide additional per-client results in Sec. A.12.
Moreover, as shown in Tab.3, not only do clients using smaller models (i.e., LLaVA-Llama3.2-1B) benefit from knowledge sharing through PFL, but larger models (i.e., LLaVA-Llama3.2-3B) also see significant gains. We attribute this to (i) SIMA accurately measuring task similarity under distribution shifts, reducing interference during aggregation, and (ii) PQ-LoRA enabling effective knowledge transfer between heterogeneous architectures, allowing smaller and larger models to mutually benefit.
We also emphasize that FedMosaic significantly outperforms baselines in generalization (‘Others’), which is crucial under distribution shifts where new, unseen tasks continuously emerge. This enhanced generalizability enables faster adaptation and accelerates future personalization, as shown in Fig. 5. The superior $A_{last}$ performance further confirms that FedMosaic effectively accumulates global knowledge during training, enabling quicker adaptation to new tasks.
Evaluation of Fast Adaptation Capability.
In real-world scenarios, unseen task data continuously emerge, making rapid adaptation crucial for future personalization. We evaluate adaptation speed by initializing models with aggregated local models from each PFL method and fine-tuning on unseen tasks for 200 iterations. As shown in Fig. 5, while sufficient training eventually leads to good downstream performance regardless of initialization, models initialized with FedMosaic achieve strong performance with just a few epochs, demonstrating that FedMosaic enhances generalizability by effectively sharing knowledge in heterogeneous PFL setups.
Heterogeneous PFL Setup in LLMs.
Beyond multi-modal PFL, we evaluate FedMosaic in heterogeneous PFL for LLM training. Tab. 4 shows FedMosaic significantly outperforms baselines in both ‘Self’ and ‘Others’, consistent with our MLLM-FL benchmark results.
Additional Experiments.
We compare accuracy and relative FLOPs in Fig. 6. As shown, FedMosaic demonstrates higher personalization performance with computation comparable to SFT.
We present experimental results for the Heterogeneous PFL setup using three heterogeneous architectures, LLaVA-Llama3.2-1B, LLaVA-Llama3.2-3B, and LLaVA-Llama3.1-8B (Sec.A.8), and LLaVA-Qwen2.5 variant models (Sec.A.10). We also present results in Homogeneous PFL setup (Sec. A.9), results in Fed-LLM benchmark (Sec. A.11), extended fast adaptation evaluations (Sec. A.17), computation analysis (Sec. A.6), details of client model selection in the Heterogeneous PFL setup(Sec. A.3).
Figure 5: Comparison of adaptation speed. We use DRAKE’s unseen tasks as downstream tasks. Random init starts from randomly initialized models, while other baselines are initialized from aggregated local models trained on DRAKE using each respective FL baseline.
Method Self Others
$A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$
SFT 2.34$\pm$0.08 2.17$\pm$0.03 2.01$\pm$0.02 1.99$\pm$0.01
FedSim (ICML 2022) 1.99$\pm$0.03 1.98$\pm$0.04 1.92$\pm$0.01 1.92$\pm$0.03
FedIT (ICASSP 2024) 2.27$\pm$0.08 2.18$\pm$0.02 2.19$\pm$0.11 2.10$\pm$0.04
TAKFL (NeurIPS 2024) 2.34$\pm$0.12 2.17$\pm$0.06 2.19$\pm$0.10 2.08$\pm$0.03
FedDPA (NeurIPS 2024) 2.39$\pm$0.09 2.27$\pm$0.02 2.20$\pm$0.05 2.13$\pm$0.01
PerAda (CVPR 2024) 2.33$\pm$0.09 2.12$\pm$0.03 2.12$\pm$0.10 1.98$\pm$0.02
FedMKT (COLING 2025) 2.31$\pm$0.06 2.12$\pm$0.05 2.06$\pm$0.02 1.96$\pm$0.00
FedMosaic (Ours) 2.51$\pm$0.01 2.32$\pm$0.03 2.25$\pm$0.01 2.17$\pm$0.02
Table 4: Quantitative comparison in heterogeneous LLM Clients on FedAya-Dynamic. We compute GPT scores evaluated by GPT-4 (Sec. A.5).
Figure 6: Accuracy and relative FLOPs in heterogeneous DRAKE-Dynamic.
6.3 Ablation Study
We ablate FedMosaic to investigate the benefit of each proposed component, and summarize the results in Tab. 5. Our observations indicate that each component plays a significant role in enhancing both the personalization (‘Self’) and the generalization (‘Others’) performance. Specifically, PQ-LoRA enhances generalization by transferring knowledge across heterogeneous architectures, while SIMA improves personalization by promoting knowledge sharing among similar-task models only.
We further present an ablation study: effect of low rank $r$ in PQ-LoRA (Sec. A.13), effect of $N_{B}$ (Sec. A.16), effect of weight alignment in PQ-LoRA (Sec. A.15), and the effect of decayed client-wise gradient $\hat{g}$ in SIMA (Sec. A.14). We discuss limitations in Sec. A.19.
Method DRAKE HFLB
Self Others Self Others
$A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$
Vanilla 66.10$\pm$0.27 57.87$\pm$0.27 47.63$\pm$0.16 46.62$\pm$0.04 79.90$\pm$0.38 77.21$\pm$0.50 60.29$\pm$1.47 60.37$\pm$1.25
(+) PQ-LoRA 66.99$\pm$0.77 58.39$\pm$0.38 51.33$\pm$0.06 48.78$\pm$0.03 80.55$\pm$0.11 77.86$\pm$0.45 66.63$\pm$0.30 65.92$\pm$0.08
(+) PQ-LoRA &SIMA (Ours) 67.86$\pm$0.51 59.83$\pm$0.16 51.16$\pm$0.04 49.36$\pm$0.08 80.90$\pm$0.30 79.18$\pm$1.04 66.66$\pm$0.77 65.80$\pm$0.30
Table 5: Ablations for proposed components of FedMosaic in heterogeneous PFL. ‘Vanilla’ aggregates homogeneous local models within each model type by averaging them with equal weights.
7 Conclusion
We address the challenges of both data heterogeneity and model heterogeneity in PFL. While prior works typically tackle only one of these challenges, or both under simplified assumptions, we propose FedMosaic, which handles both forms of heterogeneity comprising two key components: SIMA and PQ-LoRA. Moreover, to better reflect real-world data heterogeneity in PFL scenarios, we introduce DRAKE, a comprehensive multi-modal FL benchmark that captures both task heterogeneity and distribution shifts. Our extensive empirical evaluation across diverse PFL scenarios shows that FedMosaic not only achieves superior personalization, outperforming local training, but also improves generalization and few-shot adaptation, which are essential for future personalization.
Acknowledgment
We acknowledge support from the European Research Council (ERC-2020-AdG 101021347 KeepOnLearning). This work was partly supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grants funded by the Korean government (MSIT): No.RS-2022-II220077, No.RS-2022-II220113, No.RS-2022-II220959, No.RS-2022-II220871, No.RS-2022-II220951 (50%), No.RS-2021-II211343 (SNU AI), No.RS-2021-II212068 (AI Innovation Hub), and IITP-2025-RS-2024-00437633 (ITRC). Additional support was provided by the BK21 FOUR program, SNU in 2025.
Appendix A Technical Appendices and Supplementary Material
This document provides a comprehensive overview of the technical appendices
Theoretical Foundations
  • Section A.1: Theoretical justification for freezing matrices $A$ and $B$ in PQ-LoRA, proving that this approach achieves zero aggregation error in federated learning.
  • Section A.2: Proof of Theorem 1 showing that orthogonal initialization of $A$ and $B$ maximizes the representational capacity of PQ-LoRA.
Implementation Details
  • Section A.3: Details on client model selection methodology in heterogeneous PFL scenarios.
  • Section A.4: In-depth explanation of PQ-LoRA, including initialization with orthogonal sets, weight alignment, and orthogonality-enforcing post-processing.
  • Section A.5: Comprehensive description of experimental configurations, including model architecture details, benchmark specifications, evaluation metrics, and hyperparameter settings.
  • Section A.6: Comparative analysis of memory and computational costs across various baselines.
  • Section A.7: Empirical analysis supporting block-wise aggregation through CKA similarity measurements across heterogeneous models.
  • Section A.21: Detailed algorithms for the FedMosaic framework, including initialization, alignment, and training procedures.
Extended Experimental Results
  • Section A.8: Results with more diverse heterogeneous PFL scenarios, including experiments with three different model architectures.
  • Section A.9: Experimental results in homogeneous PFL setups.
  • Section A.10: Additional experiments using Qwen-based LLaVA models.
  • Section A.11: Results on the Fed-LLM benchmark, complementing our Fed-Aya experiments.
  • Section A.12: Detailed client-wise accuracy analysis for both homogeneous and heterogeneous PFL setups.
  • Section A.17: Extended evaluation of fast adaptation capabilities on additional unseen tasks.
Ablation Studies
  • Section A.13: Analysis of how the low rank parameter $r$ in PQ-LoRA affects performance.
  • Section A.14: Ablation study on SIMA, comparing different strategies for measuring task similarity.
  • Section A.15: Investigation of the effects of weight alignment in PQ-LoRA.
  • Section A.16: Study on the effect of varying the number of layers employing PQ-LoRA.
Benchmark Details and Limitations
  • Section A.18: Comprehensive description of our proposed DRAKE benchmark, including task categories, dataset sources, and client configuration.
  • Section A.19: Discussion of limitations and potential directions for future work.
  • Section A.20: Impact statement addressing the broader implications of our research.
A.1 Theoretical Justification for Freezing $A$ and $B$ in PQ-LoRA
Recap of PQ-LoRA.
PQ-LoRA consists of LoRA matrices $A\in\mathbb{R}^{r\times d_{I}}$, $B\in\mathbb{R}^{d_{O}\times r}$, along with shareable modules $P\in\mathbb{R}^{r\times r}$, $Q\in\mathbb{R}^{r}$. PQ-LoRA outputs $h_{O}$ for the given input $h_{I}$ as:
$h_{O}=W_{p}h_{I}+B(PAh_{I}+Q),$ (10)
where $W_{p}$ refers to the pre-trained weight.
For simplicity, we assume that $Q=\mathbf{0}\in\mathbb{R}^{r}$, i.e., $h_{O}=BPAh_{I}$, and prove under the homogeneous FL setup, but can easily extend to $Q\neq\mathbf{0}$ and the heterogeneous FL setup.
Definition (Aggregation Error $\delta$).
We first define the aggregation error $\delta$. In an FL setup with $N$ clients, the server aggregates client updates $\Delta W_{i}=B_{i}P_{i}A_{i}$ for $i\in[N]$. With FedAvg aggregation, the ‘ideal’ aggregation $\Delta W^{*}$ (sun2024improving, ; guo2025selective, ) should be:
$\Delta W^{*}=\frac{1}{N}(B_{1}P_{1}A_{1}+B_{2}P_{2}A_{2}+\cdots+B_{N}P_{N}A_{N}).$ (11)
However, we cannot perform ‘ideal’ aggregation in FL, since $W^{*}$ cannot be decomposed into trainable parameters, i.e., $B$, $P$, and $A$, on client sides (guo2025selective, ). In other words, although $W^{*}$ can be computed on the server, it cannot be redistributed as trainable parameters on the client side. As a result, instead of directly averaging $\Delta W_{i}$ for $i\in[N]$, we average the trainable parameters separately and obtain the practical update $\Delta\tilde{W}^{*}$ as:
$\Delta\tilde{W}^{*}=\left(\frac{1}{N}\sum_{i=1}^{N}B_{i}\right)\left(\frac{1}{N}\sum_{i=1}^{N}P_{i}\right)\left(\frac{1}{N}\sum_{i=1}^{N}A_{i}\right).$ (12)
To this end, we define the aggregation error $\delta=|\Delta W^{*}-\Delta\tilde{W}^{*}|$. During local training, we update shareable modules $P$ and $Q$, while freezing $A$ and $B$ to preserve alignment. This way, we minimize the aggregation error $\delta$, following Theorem 2.
Theorem 2.
If we freeze $A\in\mathbb{R}^{r\times d_{I}}$ and $B\in\mathbb{R}^{d_{O}\times r}$, aggregation error $\delta=0$.
Proof of Theorem 2.
Proof.
If $A$ is frozen after alignment, then $A=A_{1}=A_{2}=\cdots=A_{N}$. Similarly, freezing $B$ implies $B=B_{1}=B_{2}=\cdots=B_{N}$. Under this condition, both the ideal update $\Delta W^{*}$ (Eq. 11) and the practical update $\Delta\tilde{W}^{*}$ (Eq. 12) can be simplified to:
$\Delta W^{*}=\Delta\tilde{W}^{*}=B\left(\frac{1}{N}\sum_{i=1}^{N}P_{i}\right)A.$ (13)
Thus, $\delta=0$. ∎
A.2 Proof of Theorem 1
Proof.
For simplicity, we assume that $Q=\mathbf{0}\in\mathbb{R}^{r}$.
The weight update for PQ-LoRA is given by:
$\Delta W=BPA,$ (14)
where $B\in\mathbb{R}^{d_{O}\times r},P\in\mathbb{R}^{r\times r},A\in\mathbb{R}^{r\times d_{I}}$.
We can decompose the matrix $P$ using scalar values and basis vectors as follows:
$P=\sum_{i=1}^{r}\sum_{j=1}^{r}P_{ij}\cdot e_{i}e_{j}^{T},$ (15)
where $P_{ij}$ refers to the scalar value of $(i,j)$ position of matrix P, and $e_{i}$ and $e_{j}$ refers to the standard basis vector in $\mathbb{R}^{r}$.
Substituting Eq.15 into Eq.14, we get:
$\Delta W=BPA=B\left(\sum_{i,j}P_{ij}\cdot e_{i}e_{j}^{\text{T}}\right)A=\sum_{i,j}P_{ij}\cdot\left(Be_{i}\right)\cdot\left(e_{j}^{\text{T}}A\right).$ (16)
Since $Be_{i}=b_{i}\in\mathbb{R}^{d_{O}}$ ($i_{\text{th}}$ column of $B$), $e_{j}^{T}A=a_{j}^{\text{T}}\in\mathbb{R}^{d_{I}}$ ($j_{\text{th}}$ row of $A$), we can simplify Eq. 16 as:
$\Delta W=\sum_{i,j}P_{ij}\cdot b_{i}a_{j}^{\text{T}}.$ (17)
Therefore, representation space of $\Delta W$, denoted as $S_{PQ}$, is defined as:
$S_{PQ}=\text{span}\{b_{i}a_{j}^{\text{T}}\mid i,j=1,\dots,r\}.$ (18)
Since both $\{b_{1},b_{2},\dots,b_{r}\}\subset\mathbb{R}^{d_{O}}$ and $\{a_{1}^{\text{T}},a_{2}^{\text{T}},\dots,a_{r}^{\text{T}}\}\subset\mathbb{R}^{d_{I}}$ form orthogonal sets, their outer products $\{b_{i}a_{j}^{\text{T}}\}_{i,j=1}^{r}$ are linearly independent.
Therefore, the representation space $S_{PQ}$ has the maximum possible dimension:
$\dim(S_{PQ})=r^{2}.$ (19)
To satisfy Eq.19, it is sufficient for the column vectors of $B$ and the row vectors of $A$ to form linearly independent sets, even if they are not strictly orthogonal. However, compared to initialization with linearly independent vectors, orthogonal initialization offers additional advantages, such as faster (hu2020provable, ) and stable (nowak24a, ) convergence. Therefore, motivated by both the practical advantages and the theoretical justification provided in Theorem 1, we initialize $A$ and $B$ with orthogonal sets.
A.3 Details of Client Model Selection in the Heterogeneous PFL Scenario
In the heterogeneous PFL scenarios, we assign each client’s model based on supervised fine-tuning (SFT) performance on their respective tasks. Specifically, clients are assigned larger models when these models exhibit significantly better SFT performance for their tasks. Conversely, when the performance difference between model sizes is negligible, we allocate smaller models to maximize computational efficiency, as larger models require substantially more computational resources. We summarize the SFT performance of each LLaVA-Llama3 variant model and the resulting client-wise model allocation in Tab. 6.
Model DRAKE
Client 1 Client 2 Client 3 Client 4 Client 5 Client 6 Client 7 Client 8 Client 9 Client 10
LLaVA-Llama3.2-1B 78.43 57.98 65.05 62.53 63.15 63.35 47.56 51.68 53.07 50.95
LLaVA-Llama3.2-3B 79.33 59.01 66.75 63.33 66.00 76.15 63.40 68.21 63.08 57.32
Allocated Model LLaVA-Llama3.2-1B LLaVA-Llama3.2-3B
Model HFLB
Client 1 Client 2 Client 3 Client 4 Client 5 Client 6 Client 7 Client 8 Client 9
LLaVA-Llama3.2-1B 79.78 98.18 51,23 80.73 71.58 80.48 71.64 76.45 89.45
LLaVA-Llama3.2-3B 80.33 97.32 52.85 81.85 73.84 82.75 74.47 79.50 91.50
Allocated Model LLaVA-Llama3.2-1B LLaVA-Llama3.2-3B
Table 6: Per-client SFT performance of different LLaVA-Llama3 model variants
A.4 More Details of PQ-LoRA
As discussed in Sec.4.2.2, we enforce orthogonality on the $A$ and $B$ matrices to maximize the capacity of PQ-LoRA, following Theorem1. This is achieved through (i) initializing $A$ and $B$ with orthogonal sets, (ii) weight alignment, and (iii) orthogonality-enforcing post-processing.
Initialization with Orthogonal Set.
We initialize row vectors of $A\in\mathbb{R}^{r\times d_{I}}$ and column vectors of $B\in\mathbb{R}^{d_{O}\times r}$ with an orthogonal set, as follows:
$AA^{\top}=I_{r},\quad B^{\top}B=I_{r},$ (20)
where $I_{r}$ refers to the $r\times r$ identity matrix.
Weight Alignment.
For $B_{i}$ and $B_{j}$, we align them using canonical correlation analysis (CCA), as follows:
$B_{j}=({\Pi_{i}}^{-1})^{T}\cdot(\Pi_{i})^{T}\cdot B_{i},$ (21)
where $\Pi_{i}$ and $\Pi_{j}$ are projection matrices that project $B_{i}$ and $B_{j}$ into to the maximally correlated space, respectively. During alignment, $B_{j}$ remains orthogonal because $B_{i}$ is initialized with orthogonal vectors, and the projection in Eq. 21 preserves this orthogonality.
However, during the alignment of $A_{i}$ and $A_{j}$ in Eq. 8, orthogonality can be disrupted due to MSE loss training. To prevent significant deviation from orthogonality, we add a regularization term to the objective function of $A$ alignment as follows:
$\min_{A_{j}}\frac{1}{|\mathcal{D}_{p}|}\sum_{(x,y)\in\mathcal{D}_{p}}\mathrm{MSE}\bigl{(}A_{i}(x),A_{j}(x)\bigr{)}+\lambda\left\|A_{j}^{\top}A_{j}-I\right\|_{F}^{2},$ (22)
where $\|\cdot\|_{F}$ denotes the Frobenius norm. Despite the regularization, $A_{j}$ may not strictly satisfy orthogonality. Therefore, to ensure exact orthogonality, we perform an additional post-processing step.
Orthogonality-Enforcing Post-processing.
To enforce orthogonality in $A_{j}$, we apply orthogonal projection. Specifically, we aim to find the closest orthogonal matrix $A^{*}_{j}$ from $A_{j}$ by minimizing the Frobenius norm, as follows:
$\min_{{A^{*}_{j}}^{T}A^{*}_{j}=I}\|A_{j}-A^{*}_{j}\|_{F}.$ (23)
This optimization can be efficiently solved using the Singular Value Decomposition (SVD). We compute the SVD of $A$:
$A_{j}=U\Sigma V^{T},$ (24)
where $U\in\mathbb{R}^{r\times r}$, $\Sigma\in\mathbb{R}^{r\times d_{I}}$, and $V\in\mathbb{R}^{d_{I}\times d_{I}}$.
The closest orthogonal matrix $A^{*}_{j}$ is then given by:
$A^{*}_{j}=UV^{T}.$ (25)
As a result, we get $A_{i}$ and $B_{i}$, initialized with orthogonal sets and kept frozen during alignment, and $A^{*}_{j}$ and $B_{j}$, which are aligned with $A_{i}$ and $B_{i}$ while maintaining orthogonality.
A.5 Details of Experimental Setups
Models.
To employ LLaVA-Llama3 variant models (i.e., LLaVA-Llama3.2-1B, LLaVA-Llama3.2-3B, LLaVA-Llama3.1-8B) and LLaVA-Qwen2.5 variant models (i.e., LLaVA-Qwen2.5-0.5B, LLaVA-Qwen2.5-1.5B, and LLaVA-Qwen2.5-3B) in heterogeneous PFL scenarios, we instruction-tune them on LLaVA-Instruct-158K (liu2023visual, ), following the original LLaVA training setup (liu2023visual, ). We summarize the performance of the instruction-tuned variants in Tab. 7.
Model VQAv2 GQA VizWiz SciQA TextVQA POPE MME MMBench SEED LLaVA MM-Vet
(goyal2017vqav2, ) (hudson2019gqa, ) (gurari2018vizwiz, ) (lu2022sciqa, ) (singh2019textvqa, ) (li2023pope, ) (fu2024mme, ) (liu2024mmbench, ) -Bench (li2023seedbench, ) -Wild (liu2023visual, ) (yu2023mmvet, )
LLaVA-1.5-7B (liu2023visual, ) 78.5 62.0 50.0 66.8 58.2 85.9 1510.7 64.3 58.6 65.4 31.1
LLaVA-Llama3.2-1B 75.3 59.7 45.3 67.3 49.0 85.0 1338.1 61.3 57.8 59.6 28.7
LLaVA-Llama3.2-3B 78.7 62.6 45.3 75.71 55.8 86.1 1446.5 70.9 62.5 67.0 34.5
LLaVA-Llama3.1-8B 79.8 64.2 47.6 76.4 57.9 86.3 1486.1 72.5 64.1 73.7 32.7
LLaVA-Qwen2.5-0.5B 71.8 56.5 37.3 58.1 43.4 86.2 1251.9 53.6 53.1 55.0 23.1
LLaVA-Qwen2.5-1.5B 75.3 59.3 40.9 69.5 51.8 86.4 1376.4 66.8 60.6 60.0 27.9
LLaVA-Qwen2.5-3B 77.2 61.4 50.9 74.1 54.7 87.2 1447.6 71.4 63.2 66.7 33.1
Table 7: Zero-Shot performance of instruction-tuned LLaVA-1.5 variants. We use the CLIP ViT-L/336px model as the vision encoder for all variants.
Benchmarks.
We evaluate FedMosaic on Multi-modal FL benchmarks, such as our proposed DRAKE and HFLB, and LLM-FL benchmarks, such as Fed-LLM and Fed-Aya. We summarize the details of benchmarks in Tab. 8. As shown in the table, unlike HFLB, which focuses solely on VQA tasks, DRAKE covers diverse task types and involves more images per sample due to its multi-image inputs. In DRAKE-Dynamic, each client learns four tasks, where all data samples are randomly mixed in PFL-Static, while the samples from each task are introduced incrementally in PFL-Dynamic. We provide more detail of DRAKE in Sec. A.18.
HFLB includes 7 VQA datasets, where we assign them into 9 clients by partitioning GQA hudson2019gqa into three clients based on the QA type (Yes/No, 2 Multi-choice, and 4 Multi-choice). We further split each assigned dataset into four tasks to allow experiments on the PFL-Dynamic setup. We use keywords in questions or GPT to categorize each sample, where the details are provided in Tab. 9.
Benchmark ##\## of Images ##\## of Questions ##\## of clients ##\## of rounds $R$#\## of local steps Task types
DRAKE 357K 274K 10 20 100 Visual Relation, Multi-modal Reasoning, VQA
HFLB (chen2024feddat, ) 217K 314K 9 20 100 VQA
Fed-LLM (ye2024fedllm, ) - 30K 5 20 30 General Capability, Mathematics, Coding
Fed-Aya (singh2024aya, ) - 36K 8 20 50 Question-Answering in 12 Languages
Table 8: Benchmark details and federated learning configurations.
Client 1 Client 2 Client 3 Client 4 Client 5
GQA (hudson2019gqa, ): Yes/No
$\bullet$ Attribute
$\bullet$ Relation
$\bullet$ Object
$\bullet$ Global
GQA (hudson2019gqa, ): Four-choice
$\bullet$ Attribute
$\bullet$ Relation
$\bullet$ Category
$\bullet$ Global
GQA (hudson2019gqa, ): Two-choice
$\bullet$ Attribute
$\bullet$ Relation
$\bullet$ Category
$\bullet$ Global
Abstract VQA (antol2015abstract, )
$\bullet$ Attribute
$\bullet$ Number
$\bullet$ Yes/No
$\bullet$ Others
SNLI-VE (xie2019snlive, )
$\bullet$ Action
$\bullet$ Scene Context & Obj Relations
$\bullet$ Object-Centric
$\bullet$ Commonsense
Client 6 Client 7 Client 8 Client 9
COCOQA (ren2015cocoqa, )
$\bullet$ Object
$\bullet$ Number
$\bullet$ Color
$\bullet$ Location
NLVR2 (suhr-etal-2019nlvr, )
$\bullet$ Presupposition Negation Universal
$\bullet$ Spatial Comparative
$\bullet$ Cardinality Existential
$\bullet$ Coordination Coreference
VizWiz (gurari2018vizwiz, )
$\bullet$ Food, Brand, and Label Identification
$\bullet$ Color and Type Identification
$\bullet$ Optical Character Recognition
$\bullet$ General Object Identification
AQUA (garcia2020aqua, )
$\bullet$ “standing” keyword
$\bullet$ “sit” keyword
$\bullet$ “wear”, “hold”, “walk”, “talk” keywords
$\bullet$ Other keywords
Table 9: Per-client task configuration of HFLB. Datasets are partitioned into question-type subsets using keyword rules or GPT judgement.
Metrics.
We measured the accuracy metrics, $A_{last}$ and $A_{\text{AUC}}$, based on the correct choice for multi-choice questions and the correct tokens compared to the ground-truth answer for the open-ended questions. For Fed-Aya, we used GPT to rate the generated response compared to the given ground-truth response. We use the same prompt template shown in Fig. 7 from the original paper (singh2024aya, ). For Fed-LLM, we follow the evaluation process of the original paper (ye2024fedllm, ) and use MMLU (hendrycks2020mmlu, ), GSM8K (cobbe2021gsm8k, ), and HumanEval (chen2021humaneval, ) benchmarks to evaluate general capability, math and coding skills, respectively.
[Instruction]
Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user question displayed below. A good answer should follow these rules:
1. It should be in the same language as the question. 2. It should answer the request in the instruction. 3. It should be factually and semantically comprehensible. 4. It should be grammatically correct and fluent. Begin your evaluation by providing a short explanation. Be as objective as possible. After providing your explanation, you must rate the response on a scale of 1 to 10 by strictly following this format: “[[rating]]”, for example: “Rating: [[5]]”. A human-annotated answer is given for reference. [Question]
{question}
[The Start of Assistant’s Answer]
{answer}
[The End of Assistant’s Answer]
[Reference]
{reference}
Figure 7: Prompt template used in GPT-4 judge.
Implementation Details and Hyperparameters.
We set the batch size to 4, the learning rate for PQ-LoRA to $5\times 10^{-5}$, and for other parameters to $2\times 10^{-5}$. We use the Constant LR scheduler and the AdamW optimizer (loshchilov2019decoupled, ) for all datasets. We use the LoRA rank $r$ of 128 attached to all linear layers in the LLM backbone. For LLM experiments, we set the learning rate to $3\times 10^{-4}$ and the LoRA rank to 16. For federated learning, we set the total number of communication rounds $R$ to 20. For a fair comparison, we adjust the local training step for each baseline and FedMosaic to use the same computational cost as SFT, following  (seo2025budgeted, ). For the PFL-Dynamic setup, we employ memory-only training, where newly encountered samples are added to episodic memory, and training batches are retrieved solely from memory, following  (koh2023online, ; seo2024learning, ; seo2025budgeted, ). For episodic memory, we adopt a memory-infinite setup (prabhu2023online, ; seo2025budgeted, ), assuming all data can be stored in memory, reflecting that memory cost is not a bottleneck in real-world scenarios. For $\tau$, the softmax temperature parameter used in SIMA, we use 0.5 in all experiments. For $\lambda$, loss balancing coefficient between MSE loss and regularization term in Eq. 22, we use 0.5. We set $N_{B}$, the number of layers employing PQ-LoRA, to 4 in all experiments.
For $W_{s}$, a small-scale pre-trained MLLM used for calculating client-wise gradients in SIMA, we adopt the smallest model among all client models. For example, in Table 11, where clients have three different types of heterogeneous architectures (i.e., LLaVA-Llama3.2-1B, LLaVA-Llama3.2-3B, and LLaVA-Llama3.1-8B), we employ LLaVA-Llama3.2-1B as $W_{s}$ for all clients. This choice ensures computational efficiency and maintains gradient dimension consistency for cosine similarity calculation. All experiments are executed in Python 3.10, on four Ubuntu 20.04 machines, with 8 NVIDIA RTX A6000 GPUs each. Each experiment runs on a single RTX A6000 GPU in a day.
A.6 Comparison of Memory and Computational Costs
We compare the computational cost $\mathcal{C}$ and memory cost $\mathcal{M}$ of various baselines and summarize the results in Tab. 10. Following (seo2025budgeted, ), we measure computational cost in FLOPs per iteration and memory cost in Bytes. Specifically, for each baseline, we report the relative FLOPs and relative Bytes in comparison to supervised fine-tuning (SFT), which only requires a single forward and backward pass without any extra computation and memory overhead.
Comparison of Computational Cost $\mathcal{C}$.
We first compare the computational cost of FL methods. PerAda (xie2024perada, ) incurs approximately twice the computational cost compared to other baselines, since it sequentially updates both the personalized adapter and the local adapter. Similarly, FedDAT (chen2024feddat, ) independently optimizes the local adapter and the dual adapter teacher (DAT), which also results in a double computational cost. Our proposed FedMosaic incurs a minor additional computational cost due to gradient computation from the frozen pre-trained model used for measuring task similarity and dual adapter structure, but this overhead amounts to only approximately 9.8% more FLOPs. The reason for the small increase in computation is that we only use the last layer gradient of the small-scale pre-trained frozen model for computational efficiency, as mentioned in Sec. 4.1. Specifically, we minimize gradient computation overhead by using only the last-layer gradients of the frozen LLaVA-1.5-1B model on one batch out of every 10 iterations, rather than computing gradients on all data. Moreover, note that while FedMosaic also introduces a dual adapter structure, i.e., maintaining the local adapter and the global adapter separately, similar to PerAda, FedDAT, and FedDPA, we optimize the dual-adapter at once using a learnable balancing parameter in PQ-LoRA, which adaptively balances local and global output. In contrast, the baselines separately optimize each adapter, thus incurring a double computational cost.
Comparison of Memory Cost $\mathcal{M}$.
We then compare the memory cost of FL methods. DITTO, FedSim, PerAda, FedDAT, FedDPA, as well as FedMosaic employ a dual adapter structure, which maintains both local and global adapters separately, to preserve global knowledge. It incurs additional memory cost, but since we adopt the LoRA adapter (hu2022lora, ), which occupies significantly less memory compared to pre-trained weights (qi2024fdlora, ), the dual adapter only consumes about 5% additional memory cost compared to using a single LoRA adapter.
Methods Memory Cost $\mathcal{M}$ Computational Cost $\mathcal{C}$
Overhead Type Relative $\mathcal{M}$ to SFT Overhead Type Relative $\mathcal{C}$ to SFT
SFT - 1.000 - 1.000
DITTO (ICML 2021) Dual Adapter 1.052 Double Forward/Backward 2.000
FedSim (ICML 2022) Dual Adapter 1.052 Double Forward 1.487
FedIT (ICASSP 2024) - 1.000 - 1.000
TAKFL (NeurIPS 2024) - 1.000 Distill Logit Extract 1.294
PerAda (CVPR 2024) Dual Adapter 1.052 Double Forward/Backward & Logit Extraction 2.294
FedDAT (AAAI 2024) Triple Adapter 1.104 Double Forward/Backward 2.564
FedDPA (NeurIPS 2024) Dual Adapter 1.052 Double Forward/Backward 2.051
FedMKT (COLING 2025) - 1.000 Logit Extraction 1.574
FedMosaic (Ours) PQ-LoRA 1.053 Last Layer Gradient Compute 1.098
Table 10: Comparison of memory and computational costs. FedMosaic incurs additional computational and memory costs compared to SFT, but only by approximately 5.3% and 9.8%, respectively.
A.7 Empirical Analysis of Block-Wise Aggregation
To identify layer-wise correspondences between depth-heterogeneous models, we analyze representation alignment using CKA (kornblith2019similarity, ). Specifically, we measure similarity across layers within the LLaMA-3 family (1B, 3B, 8B) and the Qwen-2.5 family (0.5B, 1.5B, 3B), as illustrated in Fig. 8. As shown in the figure, layers with the same relative depth exhibit strong alignment, indicating approximately linear alignment within both the LLaMA-3 and Qwen-2.5 families. Moreover, we observe near-linear alignment even across families, i.e., between LLaMA-3 and Qwen-2.5, despite weaker linearity than intra-family alignment. This empirical analysis supports our block-wise aggregation of PQ-LoRA. We provide an illustration of the block-wise PQ-LoRA in Fig. 9.
(a) Llama-3 family (1 B, 3 B, 8 B)
(b) Qwen-2.5 family (0.5 B, 1.5 B, 3 B)
(c) Llama - Qwen mixed
Figure 8: Layer-wise similarity across model scales measured with CKA. Each heat-map cell reports the centered-kernel–alignment (CKA) similarity between the hidden representations of heterogeneous multi-modal LLMs at every pair of layers. Here, lighter colors indicate higher similarity. (a) For the Llama-3–based heterogeneous models and (b) for the Qwen-2.5–based heterogeneous models, the brightest (i.e., the highest similarity) band appears roughly along the main diagonal, indicating that layers with relative depth align most strongly. The near-linear trend supports our proposed block-wise aggregation, which transfers knowledge from smaller to larger models that have the same relative depth within the same architectural family.
Figure 9: Illustration of blockwise PQ-LoRA. When a model has $N_{B}$ PQ-LoRA modules, each block employs PQ-LoRA at its last layer, while the remaining layers adopt conventional LoRA. Each block contains the same number of layers.
A.8 Experimental Results in More Diverse Heterogeneous PFL Scenarios
We evaluate FedMosaic in heterogeneous PFL using three different heterogeneous architectures, i.e., LLaVA-Llama3.2-1B, LLaVA-Llama3.2-3B, and LLaVA-Llama3.1-8B, and summarize the results in Tab. 11. Consistent with the previous heterogeneous PFL scenario with two different types of architectures (i.e., LLaVA-Llama3.2-1B and LLaVA-Llama3.2-3B), FedMosaic consistently outperforms the baselines in both PFL-Dynamic and PFL-Static settings. These results demonstrate that PQ-LoRA effectively enables knowledge sharing across various heterogeneous architectures, highlighting its applicability to real-world scenarios where each client possesses individual heterogeneous models.
Method DRAKE-Dynamic DRAKE-Static
Self Others Self Others
$A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$
SFT 66.41$\pm$0.84 59.17$\pm$0.64 47.94$\pm$0.18 47.10$\pm$0.19 68.86$\pm$2.16 62.30$\pm$3.12 47.98$\pm$0.27 47.44$\pm$0.20
DITTO (ICML 2021) 61.56$\pm$0.01 56.36$\pm$0.20 48.19$\pm$0.26 47.41$\pm$0.01 65.48$\pm$0.97 60.34$\pm$0.32 48.49$\pm$0.03 48.09$\pm$0.01
FedSim (ICML 2022) 65.00$\pm$0.41 58.18$\pm$0.46 47.47$\pm$0.10 46.61$\pm$0.04 67.39$\pm$0.32 62.42$\pm$0.25 47.08$\pm$0.05 47.10$\pm$0.03
FedIT (ICASSP 2024) 66.18$\pm$0.22 58.95$\pm$0.02 47.54$\pm$0.02 46.97$\pm$0.01 68.78$\pm$0.22 64.07$\pm$0.92 47.34$\pm$0.55 47.33$\pm$0.29
TAKFL (NeurIPS 2024) 64.73$\pm$0.49 58.33$\pm$0.53 47.85$\pm$0.46 47.16$\pm$0.08 68.17$\pm$0.75 63.19$\pm$1.21 47.28$\pm$0.14 46.90$\pm$1.00
FedDPA (NeurIPS 2024) 63.26$\pm$0.09 57.31$\pm$0.02 48.04$\pm$0.26 47.21$\pm$0.07 67.71$\pm$0.95 63.03$\pm$0.24 48.18$\pm$0.53 48.22$\pm$0.13
FedDAT (AAAI 2024) 60.43$\pm$0.54 56.82$\pm$0.29 49.79$\pm$0.21 48.25$\pm$0.25 61.97$\pm$0.58 57.64$\pm$0.45 49.76$\pm$0.46 47.99$\pm$0.21
PerAda (CVPR 2024) 58.10$\pm$2.74 54.84$\pm$1.68 46.90$\pm$0.22 46.69$\pm$0.21 61.57$\pm$0.97 56.22$\pm$0.38 45.01$\pm$0.37 47.06$\pm$0.71
FedMKT (COLING 2025) 63.25$\pm$0.05 57.23$\pm$0.18 47.88$\pm$0.29 47.25$\pm$0.15 65.50$\pm$0.37 61.12$\pm$0.56 48.27$\pm$0.04 47.70$\pm$0.31
FedMosaic (Ours) 68.94$\pm$0.68 60.96$\pm$0.06 52.18$\pm$0.34 50.11$\pm$0.03 70.41$\pm$1.27 65.12$\pm$1.15 52.67$\pm$0.20 50.91$\pm$0.70
Table 11: Quantitative comparison in heterogeneous PFL with three different types of models. ‘Self’ denotes evaluation on a client’s own data, while ’Others’ denotes evaluation on data from other clients. 3 clients use LLaVA-Llama3.2-1B model, 5 clients use LLaVA-Llama3.2-3B model, and 2 clients use LLaVA-Llama3.1-8B model. SFT refers to supervised fine-tuning on each client’s data without cross-client knowledge sharing.
A.9 Experiment Results on Homogeneous PFL Setup
In addition to the heterogeneous PFL scenario (Sec. 6.2), where both data distributions and model architectures vary across clients, we also evaluate FedMosaic in a homogeneous PFL setting, where clients share the same model architecture but have heterogeneous data distributions. We assume all clients use LLaVA-3B models and summarize the results for the PFL-Dynamic scenario in Tab.13 and for the PFL scenario in Tab.13. Consistent with results in the heterogeneous setup, FedMosaic significantly outperforms baselines in generalization ability (i.e., ‘Others’) and even surpasses SFT in personalization performance (i.e., ‘Self’).
Method DRAKE-Dynamic HFLB-Dynamic
Self Others Self Others
$A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$
SFT 67.57$\pm$0.14 59.18$\pm$1.26 49.70$\pm$0.02 48.04$\pm$1.18 80.26$\pm$1.14 77.98$\pm$0.49 63.24$\pm$0.23 62.90$\pm$0.60
DITTO (ICML 2021) 64.09$\pm$0.55 57.22$\pm$0.01 49.92$\pm$0.27 49.05$\pm$0.07 79.34$\pm$0.41 76.72$\pm$0.84 64.10$\pm$0.40 63.57$\pm$0.06
FedSim (ICML 2022) 67.05$\pm$0.98 60.69$\pm$0.84 47.94$\pm$1.69 47.55$\pm$1.14 79.60$\pm$0.77 76.93$\pm$0.78 60.85$\pm$1.15 60.50$\pm$0.82
FedIT (ICASSP 2024) 67.62$\pm$0.49 60.35$\pm$0.69 49.84$\pm$0.19 48.95$\pm$0.13 79.34$\pm$0.59 78.40$\pm$1.21 62.51$\pm$1.30 62.38$\pm$0.96
TAKFL (NeurIPS 2024) 67.02$\pm$0.04 59.86$\pm$0.36 50.03$\pm$0.06 48.85$\pm$0.09 79.45$\pm$0.69 77.15$\pm$0.99 63.01$\pm$0.66 62.70$\pm$0.96
FedDPA (NeurIPS 2024) 66.59$\pm$0.86 58.68$\pm$0.71 49.92$\pm$0.24 48.81$\pm$0.13 80.57$\pm$0.62 77.65$\pm$0.61 63.58$\pm$0.30 63.02$\pm$0.09
FedDAT (AAAI 2024) 61.95$\pm$0.90 57.22$\pm$0.45 51.50$\pm$0.13 50.28$\pm$0.08 79.58$\pm$0.53 76.82$\pm$1.13 67.33$\pm$0.13 66.47$\pm$0.29
PerAda (CVPR 2024) 63.78$\pm$0.70 57.44$\pm$0.16 49.79$\pm$0.25 49.10$\pm$0.02 78.88$\pm$0.96 76.31$\pm$1.24 63.35$\pm$1.26 63.09$\pm$0.56
FedMKT (COLING 2025) 65.20$\pm$0.43 58.73$\pm$0.12 49.68$\pm$0.29 48.84$\pm$0.22 79.43$\pm$0.69 77.24$\pm$0.61 62.66$\pm$0.38 62.79$\pm$0.86
FedMosaic (Ours) 70.80$\pm$0.23 62.18$\pm$0.64 53.93$\pm$0.36 51.72$\pm$0.28 80.75$\pm$0.14 78.87$\pm$0.20 68.30$\pm$0.69 67.46$\pm$0.22
Method DRAKE-Static HFLB-Static
Self Others Self Others
$A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$
SFT 70.56$\pm$0.27 66.31$\pm$0.37 50.37$\pm$0.04 49.99$\pm$0.16 80.89$\pm$0.85 80.01$\pm$0.12 63.40$\pm$0.41 63.09$\pm$0.17
DITTO (ICML 2021) 66.19$\pm$0.38 62.05$\pm$0.69 50.59$\pm$0.05 49.95$\pm$0.08 80.37$\pm$0.37 79.05$\pm$0.25 63.93$\pm$0.24 63.52$\pm$0.10
FedSim (ICML 2022) 68.65$\pm$0.86 63.96$\pm$0.78 49.27$\pm$0.22 48.85$\pm$0.14 80.05$\pm$0.01 79.36$\pm$0.79 61.22$\pm$0.30 61.31$\pm$0.12
FedIT (ICASSP 2024) 70.14$\pm$0.03 66.08$\pm$0.59 50.09$\pm$0.15 49.87$\pm$0.26 80.63$\pm$1.23 79.97$\pm$0.01 63.34$\pm$0.11 63.09$\pm$0.04
TAKFL (NeurIPS 2024) 68.75$\pm$0.04 64.39$\pm$0.07 49.70$\pm$0.16 49.60$\pm$0.03 79.97$\pm$0.11 79.41$\pm$0.81 63.67$\pm$0.39 63.30$\pm$0.25
FedDPA (NeurIPS 2024) 67.40$\pm$2.73 63.71$\pm$1.24 50.45$\pm$0.38 49.92$\pm$0.30 79.99$\pm$1.23 79.02$\pm$0.35 63.45$\pm$0.07 63.11$\pm$0.02
FedDAT (AAAI 2024) 64.09$\pm$1.16 61.10$\pm$0.60 52.20$\pm$0.26 51.47$\pm$0.04 79.80$\pm$0.50 78.43$\pm$0.68 68.92$\pm$0.44 67.66$\pm$0.27
PerAda (CVPR 2024) 65.64$\pm$0.48 61.93$\pm$0.58 50.53$\pm$0.02 49.94$\pm$0.05 80.25$\pm$0.45 79.01$\pm$0.26 64.00$\pm$0.13 63.58$\pm$0.31
FedMKT (COLING 2025) 67.85$\pm$0.92 62.93$\pm$1.16 49.23$\pm$1.54 48.77$\pm$1.69 79.85$\pm$0.67 79.72$\pm$0.60 63.84$\pm$0.20 63.39$\pm$0.07
FedMosaic (Ours) 72.02$\pm$0.37 67.18$\pm$0.13 54.21$\pm$0.05 52.94$\pm$0.05 81.69$\pm$0.16 80.54$\pm$0.09 68.70$\pm$0.26 67.75$\pm$0.23
Table 12: Quantitative comparison in homogeneous PFL. ‘Self’ denotes evaluation on a client’s own data, while ’Others’ denotes evaluation on data from other clients. All clients use LLaVA-3B models. SFT refers to supervised fine-tuning on each client’s data without cross-client knowledge sharing.
A.10 Experiment Results on Qwen-based LLaVA
In addition to heterogeneous PFL scenarios with LLaVA-Llama3 variants, we also compare FedMosaic with baselines using Qwen-based LLaVA models. Specifically, we use LLaVA-Qwen2.5-0.5B for the small model and LLaVA-Qwen2.5-3B for the large model. For baselines, we select the top-3 well-performing baselines in heterogeneous PFL scenarios using LLaVA-Llama3 variants, as well as supervised fine-tuning (i.e., SFT). We summarize the results in Tab. 13.
As shown in the table, FedMosaic significantly outperforms baselines in both homogeneous and heterogeneous PFL scenarios, consistent with results from LLaVA-Llama3 variants. This demonstrates that our proposed PQ-LoRA consistently facilitates knowledge sharing across heterogeneous models and SIMA reduces interference during local model aggregation, regardless of architecture.
Method DRAKE-Homonegeous DRAKE-Heterogeneous
Self Others Self Others
$A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$
SFT 68.79$\pm$1.36 59.06$\pm$0.18 47.87$\pm$0.40 47.10$\pm$0.04 65.79$\pm$0.17 58.24$\pm$0.14 44.40$\pm$0.86 44.06$\pm$0.43
FedIT (ICASSP 2024) 68.40$\pm$0.07 60.58$\pm$0.41 49.70$\pm$0.52 48.54$\pm$0.07 66.02$\pm$0.72 58.26$\pm$1.06 44.47$\pm$0.30 44.06$\pm$0.43
FedDPA (NeurIPS 2024) 66.73$\pm$0.42 60.21$\pm$0.47 49.39$\pm$0.07 48.69$\pm$0.11 65.43$\pm$0.64 57.79$\pm$0.08 44.48$\pm$0.43 44.53$\pm$0.26
FedMKT (COLING 2025) 65.95$\pm$0.35 58.81$\pm$0.64 49.73$\pm$0.09 48.52$\pm$0.24 62.85$\pm$0.99 56.71$\pm$0.33 44.73$\pm$1.35 44.24$\pm$0.75
FedMosaic (Ours) 70.38$\pm$0.34 62.84$\pm$0.23 54.25$\pm$0.33 51.85$\pm$0.20 67.36$\pm$0.21 59.95$\pm$0.71 50.43$\pm$0.01 48.51$\pm$0.29
Table 13: Quantitative comparison in PFL-Dynamic using Qwen-based LLaVA. ‘Self’ denotes evaluation on a client’s own data, while ‘Others’ denotes evaluation on data from other clients. In DRAKE-Homogeneous, all clients use LLaVA-Qwen2.5-3B models, while in DRAKE-Heterogeneous, 3 clients use LLaVA-Qwen2.5-0.5B model, 2 clients use LLaVA-Qwen2.5-1.5B model, and 5 client uses LLaVA-Qwen2.5-3B model. SFT refers to supervised fine-tuning on each client’s data without cross-client knowledge sharing.
A.11 Experimental Results on Fed-LLM
In addition to Fed-Aya (Sec.6.2), we evaluate FedMosaic on another LLM-FL benchmark, Fed-LLM, and summarize the results in Tab. 14. As shown in the table, FedMosaic outperforms baselines in both personalization and generalization, consistent with results from other MLLM-FL benchmarks, such as DRAKE and HFLB.
Method Self Others
$A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$
SFT 25.21$\pm$1.31 29.35$\pm$0.61 25.83$\pm$1.03 27.30$\pm$0.73
FedIT (ICASSP 2024) 24.06$\pm$0.83 28.42$\pm$0.45 26.67$\pm$0.66 28.28$\pm$0.38
FedDPA (NeurIPS 2024) 24.93$\pm$1.06 27.51$\pm$0.55 26.72$\pm$0.67 27.69$\pm$0.27
FedMKT (COLING 2025) 24.92$\pm$0.96 27.51$\pm$0.63 29.55$\pm$0.43 30.57$\pm$0.19
FedMosaic (Ours) 31.24$\pm$0.13 33.84$\pm$0.53 29.26$\pm$0.33 31.70$\pm$0.19
Table 14: Quantitative comparison of heterogeneous LLM clients on the FedLLM-Static. Three clients use the LLaMA-3B model, while the remaining two use the LLaMA-8B model.
A.12 Client-Wise Accuracy
In addition to reporting the average accuracy across all clients in Sec. 6.2, we report client-wise accuracy in both homogeneous and heterogeneous PFL setups.
Homogeneous PFL Setups.
We first compare client-wise performance in the homogeneous PFL-Dynamic setups on DRAKE and HFLB, and summarize the results in Tab.15 and Tab.16, respectively. As shown in the tables, personalization performance is improved for all clients compared to SFT (i.e., supervised fine-tuning on local data) except Client 8 in Tab. 16. We highlight this result because SFT is a strong baseline for personalization (ghari2024personalized, ), as we also show in Tab. 13, where it outperforms all baselines except for FedMosaic. Moreover, supervised fine-tuning may be sufficient for personalization in some tasks (wozniak2024personalized, ; mosbach2021on, ), as seen with Client 6 in Tab. 15, where there are only marginal improvements compared to SFT. However, for more challenging tasks, properly leveraging knowledge from other clients significantly improves personalization, e.g., FedMosaic shows a gain of 8.2% in $A_{last}$ for Client 9 and 6.4% for Client 7, as well as a 7.1% improvement in $A_{\text{AUC}}$ for Client 5. In summary, even though there is only a 1–3% gain in the average performance across clients, this is because in clients where SFT is sufficient for personalization, the gain from PFL seems smaller. Consistent with the PFL-Dynamic setup, FedMosaic outperforms SFT in the PFL-Static scenario, as shown in Tab. 17.
Method $A_{last}$ / $A_{\text{AUC}}$
Client 1 Client 2 Client 3 Client 4 Client 5 Client 6 Client 7 Client 8 Client 9 Client 10
SFT 81.90 / 62.11 65.23 / 59.28 67.38 / 60.31 70.28 / 60.40 63.69 / 62.07 76.05 / 67.35 63.40 / 54.17 66.93 / 59.31 62.70 / 52.20 58.16 / 54.58
FedMosaic (Ours) 82.73 / 66.11 67.31 / 63.76 68.97 / 62.53 72.60 / 67.51 68.67 / 63.58 76.98 / 68.56 69.83 / 58.86 70.11 / 60.08 70.89 / 54.89 59.88 / 55.89
Difference +0.83 / +4.01 +4.98 / +1.51 +2.08 / +4.48 +1.59 / +2.23 +2.31 / +7.12 +0.93 / +1.21 +6.43 / +4.69 +3.18 / +0.77 +8.19 / +2.69 +1.71 / +1.31
Table 15: Per-client performance of FedMosaic vs. SFT in a homogeneous PFL-Dynamic setup on DRAKE. The Difference row shows the performance gain/loss of FedMosaic compared to SFT. All clients use the LLaVA-Llama3.2-3B model.
Method $A_{last}$ / $A_{\text{AUC}}$
Client 1 Client 2 Client 3 Client 4 Client 5 Client 6 Client 7 Client 8 Client 9
SFT 82.28 / 79.13 96.26 / 96.59 52.73 / 50.22 82.16 / 81.52 72.79 / 72.93 86.65 / 84.74 76.04 / 73.79 81.13 / 79.83 90.58 / 83.02
FedMosaic (Ours) 82.77 / 79.79 96.60 / 97.08 52.99 / 50.88 81.89 / 81/98 74.31 / 74.95 88.20 / 86.55 76.78 / 74.74 81.94 / 80.33 90.00 / 83.25
Difference +0.49 / +0.67 +0.34 / +0.48 +0.26 / +0.67 -0.27 / +0.46 +1.52 / +2.02 +1.55 / +1.81 +0.74 / +0.94 +0.81 / +0.50 -0.58 / +0.23
Table 16: Per-client performance of FedMosaic vs. SFT in a Homogeneous PFL-Dynamic setup on HFLB. The Difference row shows the performance gain/loss of FedMosaic compared to SFT. All clients use the LLaVA-Llama3.2-3B model.
Method $A_{last}$ / $A_{\text{AUC}}$
Client 1 Client 2 Client 3 Client 4 Client 5 Client 6 Client 7 Client 8 Client 9
SFT 83.09 / 81.54 96.62 / 96.70 51.81 / 51.38 80.62 / 80.32 72.34 / 73.66 92.21 / 90.82 77.56 / 75.12 82.66 / 80.74 91.04 / 89.83
FedMosaic (Ours) 83.21 / 81.98 97.96 / 97.68 53.29 / 52.24 81.59 / 81.92 76.20 / 74.54 92.34 / 91.13 77.77 / 75.59 81.80 / 80.15 91.08 / 89.99
Difference +0.12 / +0.44 +1.33 / +0.98 +1.48 / +0.86 +0.97 / +1.60 +3.86 / +0.87 +0.13 / +0.31 +0.21 / +0.46 -0.86 / -0.59 +0.05 / +0.16
Table 17: Per-client performance of FedMosaic vs. SFT in a homogeneous PFL-Static setup on HFLB. The Difference row shows the performance gain/loss of FedMosaic compared to SFT. All clients use the LLaVA-Llama3.2-3B model.
Heterogeneous PFL Setups.
Consistent with the results in the homogeneous PFL setups, FedMosaic enhances personalization ability in most clients, as shown in Tab.18 and Tab.19. As shown in the tables, not only do clients using smaller models (i.e., LLaVA-Llama3.2-1B, LLaVA-Llama3.2-3B) benefit from knowledge sharing through PFL, but clients with larger models (i.e., LLaVA-Llama3.2-8B) also see significant improvements, e.g., a 6.5% improvement in $A_{last}$ and 9.4% improvement in $A_{\text{AUC}}$ for Client 9 in Tab. 18. We believe this is due to PQ-LoRA effectively sharing knowledge between heterogeneous architectures, allowing both smaller and larger models to provide meaningful information to each other.
Method $A_{last}$ / $A_{\text{AUC}}$
LLaVA-Llama3.2-1B LLaVA-Llama3.2-3B LLaVA-Llama3.1-8B
Client 1 Client 2 Client 3 Client 4 Client 5 Client 6 Client 7 Client 8 Client 9 Client 10
SFT 77.31 / 66.56 61.22 / 57.80 67.59 / 62.22 77.84 / 71.02 66.82 / 57.36 66.56 / 56.91 66.49 / 64.58 62.06 / 58.61 64.63 / 57.17 72.05 / 69.91
FedMosaic (Ours) 78.26 / 67.94 61.35 / 59.93 67.61 / 63.08 78.56 / 74.71 72.35 / 60.50 71.45 / 62.36 67.90 / 65.63 61.88 / 59.48 71.12 / 66.58 73.65 / 70.97
Difference +0.95 / +1.37 +0.13 / +2.13 +0.02 / +0.85 +0.73 / +3.70 +5.53 / +3.14 +4.89 / +5.45 +1.41 / +1.05 -0.19 / +0.87 +6.49 / +9.41 +1.60 / +1.05
Table 18: Per-client performance of FedMosaic vs. SFT in a heterogeneous PFL-Static setup on DRAKE. The Difference row shows the performance gain/loss of FedMosaic compared to SFT. Clients 1-3 use LLaVA-Llama3.2-1B, Clients 4-8 use LLaVA-Llama3.2-3B, and Clients 9-10 use LLaVA-Llama3.1-8B.
Method $A_{last}$ / $A_{\text{AUC}}$
LLaVA-Llama3.2-1B LLaVA-Llama3.2-3B
Client 1 Client 2 Client 3 Client 4 Client 5 Client 6 Client 7 Client 8 Client 9
SFT 80.69 / 77.18 97.41 / 95.25 50.95 / 47.72 82.39 / 81.49 73.99 / 73.00 92.10 / 88.23 76.58 / 74.19 81.00 / 79.94 89.32 / 83.12
FedMosaic (Ours) 81.69 / 77.53 96.46 / 96.18 51.38 / 47.85 82.70 / 82.41 74.49 / 75.09 91.80 / 88.36 76.95 / 74.92 82.24 / 80.25 89.94 / 83.62
Difference +1.00 / +0.36 -0.96 / +0.93 +0.43 / +0.13 +0.32 / +0.92 +0.49 / +2.08 -0.30 / +0.13 +0.36 / +0.73 +1.24 / +0.31 +0.62 / +0.50
Table 19: Per-client performance of FedMosaic vs. SFT in a heterogeneous PFL-Dynamic setup on HFLB. The Difference row shows the performance gain/loss of FedMosaic compared to SFT. Clients 1-3 use LLaVA-Llama3.2-1B, while Clients 4-9 use LLaVA-Llama3.2-3B.
A.13 Effect of Low Rank $r$ in PQ-LoRA
We study how varying the rank $r$ in PQ-LoRA affects performance, and summarize the results in Fig. 10. As shown in the figure, while both excessively high and low values of $r$ lead to degraded performance, a broad range of intermediate values shows stable performance. While a larger $r$ increases shareable capacity ($\mathbb{R}^{r\times r}$), it may cause overfitting (yang2024lora, ; cho2024heterogeneous, ). In contrast, a smaller $r$ helps mitigate overfitting but may introduce capacity limitations, resulting in suboptimal performance (he2022towards, ). By balancing the trade-off, we select an adequate low rank $r$.
Figure 10: Accuracy on different low rank $r$ values in PQ-LoRA in PFL-Dynamic setup on DRAKE.
A.14 Ablation Study of SIMA
Our proposed SIMA measures task similarity among clients using client-wise gradients to construct a similarity-aware customized global model for each client. Specifically, for the $i$-th client, we maintain a decayed gradient $\hat{g}_{i}$, updated via exponential moving average (EMA) from the current gradient $g_{i}$, which is computed using the current batch and a small, frozen pre-trained model. Note that we use the last-layer gradient of the frozen model, not the training model, as mentioned in Sec. 4.1.
We ablate SIMA by comparing four aggregation variants: (1) Equal-weight aggregation, which ignores task similarity and aggregates local models uniformly; (2) Training model gradients, using current batch gradients from the trainable model; (3) Frozen model gradients, using current batch gradients from the frozen pre-trained model; and (4) Decayed frozen gradient (i.e., SIMA), which maintains an EMA of gradients from the frozen pre-trained model to capture model knowledge shifts. We summarize the results in Tab. 20.
As shown in the table, model aggregation considering task similarity measured by client-wise gradients generally improves the ‘Self’ accuracy. This demonstrates the loss of information in Equal-weight aggregation due to parameter interference. While using gradients from the training model (‘Training model gradient’) shows improved performance compared to Equal-weight aggregation, it still suffers from data heterogeneity, as gradient similarities from models trained on heterogeneous data may not capture actual similarity (tang2020similarity, ; evans2024data, ). Using a frozen model’s gradients from the current batch (‘Frozen model gradients’) can address this limitation, but cannot reflect model knowledge shifts under distribution shifts, as mentioned in Sec. 4.1. In contrast, SIMA effectively captures task similarity under data heterogeneity with distribution shifts, thus outperforming other aggregation strategies.
Method Self
$A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$
Equal-weight aggregation 66.99$\pm$0.77 58.39$\pm$0.38
Training model gradients 67.16$\pm$0.74 59.38$\pm$0.53
Frozen model gradients 67.24$\pm$1.27 59.45$\pm$0.74
Decayed frozen gradients (SIMA) 67.86$\pm$0.51 59.83$\pm$0.16
Table 20: Ablation of components in SIMA in Heterogeneous PFL-dynamic setup on DRAKE. ‘Equal-weight aggregation’ refers to aggregating local models with equal weight, ‘Training model gradients’ refers to using gradients from the current batch computed with the training model, ‘Frozen model gradients’ refers to using gradients from the current batch computed with the frozen pre-trained model, and ‘Decayed frozen gradients’ refer to EMA-estimated gradients from the frozen model (i.e., SIMA).
A.15 Different Weight Alignment Methods in PQ-LoRA
To ensure shared initialization across heterogeneous architectures, we align $A$ matrices from heterogeneous architectures (i.e., $A_{i}\in\mathbb{R}^{r\times d_{i}}$, $A_{j}\in\mathbb{R}^{r\times d_{j}}$) using MSE loss, and $B$ matrices (i.e., $B_{i}\in\mathbb{R}^{d^{\prime}_{i}\times r}$, $B_{j}\in\mathbb{R}^{d^{\prime}_{j}\times r}$) using canonical correlation analysis (CCA), as detailed in Sec.4.2.2. We ablate the effects of aligning $A$ and $B$ matrices, and summarize the results in Tab. 21. Note that the results in the table report the average performance across all clients, which can make the overall average improvements appear small, as mentioned in Sec. 6.2.
As shown, alignment improves both personalization (‘Self’) and generalization (‘Others’) performance. This occurs because aligning both $A$ and $B$ in PQ-LoRA ensures heterogeneous models share initialization and follow consistent optimization paths, allowing aggregation without weight interference (wortsman2022model, ; wortsman2022robust, ; yadav2023ties, ). In contrast, misalignment in either matrix introduces initialization discrepancies between models, leading to weight interference during aggregation and degrading performance (jordanr2023epair, ; neyshabur2020being, ; stoica2025model, ).
Initialization Self Others
A B $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$
Random Random 67.95$\pm$1.71 63.50$\pm$0.66 47.71$\pm$0.08 47.46$\pm$0.05
Random Aligned (CCA) 68.22$\pm$1.84 63.59$\pm$0.62 47.80$\pm$0.06 47.56$\pm$0.08
Aligned (MSE) Random 68.29$\pm$1.57 63.58$\pm$0.60 47.72$\pm$0.14 47.41$\pm$0.04
Aligned (MSE) Aligned (CCA) 68.73$\pm$1.92 63.82$\pm$0.59 48.04$\pm$0.04 47.57$\pm$0.08
Table 21: Accuracy on different weight alignment in PQ-LoRA under Heterogeneous PFL-Static setup on DRAKE. ‘Random’ initializes matrices randomly, while Aligned (MSE) and Aligned (CCA) use matrices aligned by MSE loss and canonical correlation analysis (CCA), respectively.
A.16 Effect of $N_{B}$ in PQ-LoRA
We ablate the effect of $N_{B}$, the number of layers employing PQ-LoRA, and summarize the results in Tab.22. In a model $W$ with $|W|$ layers, increasing $N_{B}$ reduces the layers using conventional LoRA to $|W|-N_{B}$. Since PQ-LoRA has fewer trainable parameters (i.e., $r^{2}+r$) than conventional LoRA (i.e., $r\times(d_{I}+d_{O})$), increasing $N_{B}$ decreases the total number of learnable parameters in $W$, as $A\in\mathbb{R}^{r\times d_{I}}$ and $B\in\mathbb{R}^{d_{O}\times r}$ are frozen, with only $P\in\mathbb{R}^{r\times r}$ and $Q\in\mathbb{R}^{r}$ being trainable. Despite the reduced number of parameters, as shown in the SFT performance in the table, performance remains stable even as $N_{B}$ increases. We attribute this to orthogonal and frozen $A$ and $B$, which maximize and preserve the expressiveness of PQ-LoRA (Theorem1), despite the lower trainable parameter count.
However, there is a trade-off in $N_{B}$ in the PFL setup: increasing $N_{B}$ improves generalizability (e.g., $N_{B}=2$ vs. $N_{B}=4$ in $A_{\text{AUC}}$ of ‘Others’ in FedMosaic) but can degrade personalization (e.g., $N_{B}=4$ vs. $N_{B}=8$ in $A_{\text{last}}$ and $A_{\text{AUC}}$ of ‘Self’ in FedMosaic). This is because, while increasing $N_{B}$ enables heterogeneous models to share knowledge across more layers, the number of trainable parameters for local training decreases due to the reduction in the number of conventional LoRA modules. As a result, by balancing this trade-off, we employ a moderate value of $N_{B}$, i.e., $N_{B}=4$.
Method Self Others
$A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$
SFT ($N_{B}=2$) 66.14$\pm$0.02 57.02$\pm$0.09 47.75$\pm$0.42 46.02$\pm$0.14
SFT ($N_{B}=4$) 66.24$\pm$0.68 57.71$\pm$0.27 47.63$\pm$0.15 46.62$\pm$0.14
SFT ($N_{B}=8$) 66.21$\pm$0.62 57.87$\pm$0.20 47.86$\pm$0.29 46.76$\pm$0.20
FedMosaic ($N_{B}=2$) 68.07$\pm$0.10 59.90$\pm$0.68 51.07$\pm$0.06 48.78$\pm$0.15
FedMosaic ($N_{B}=4$) 67.86$\pm$0.51 59.83$\pm$0.16 51.26$\pm$0.04 49.36$\pm$0.08
FedMosaic ($N_{B}=8$) 67.52$\pm$0.77 59.39$\pm$0.23 51.27$\pm$0.27 49.26$\pm$0.14
Table 22: Accuracy on different number of blocks ($N_{B}$) in PQ-LoRA in heterogeneous PFL-dynamic setup on DRAKE. Given a model $W$ with $|W|$ layers, using $N_{B}$ blocks means that $N_{B}$ layers adopt PQ-LoRA, while the remaining $|W|-N_{B}$ layers use conventional LoRA.
A.17 Extended Fast Adaptation Evaluations
In addition to the fast adaptation evaluation in Sec.6.2, we further assess fast adaptation on additional unseen tasks from DRAKE. We summarize the results in Fig.11. As shown in the figure, models initialized with FedMosaic adapt significantly faster than those with random initialization or other PFL baselines, demonstrating that FedMosaic enhances generalizability by effectively sharing knowledge in heterogeneous PFL setups.
Figure 11: Comparison of adaptation speed. We use FedMosaic’s unseen tasks as downstream tasks. Random init starts from randomly initialized models, while other baselines are initialized from aggregated local models trained on DRAKE using each respective FL baseline.
A.18 Additional Details of DRAKE
We curate DRAKE using open-sourced datasets from the Internet. Specifically, we select large-scale multi-modal benchmarks that differ in distribution or properties from LLaVA’s pre-training data and provide clear metadata or instructions, enabling task-wise splits for emulating distribution shifts. We refer to diverse multi-modal benchmarks, such as DEMON (li2024demon, ), SEED-Bench-2 (li2024seed2, ), Co-Instruct (wu2024coinstruct, ), and HFLB (chen2024feddat, ).
DRAKE consists of 40 distinct tasks, where for each task, we subsample 10,000 training and 1,000 test samples, if the dataset is large; otherwise, we split the data into training and test sets with an 8:2 ratio, totaling 375k images and 274k questions. We illustrate the overall structure of DRAKE as a tree diagram in Fig. 12. As shown in the figure, DRAKE consists of three sub-groups, i.e., visual relation, multi-modal reasoning, and VQA, each comprising multiple fine-grained tasks, with representative examples of each sub-group in Fig. 13. We also provide detailed per-client task configuration of DRAKE in Tab. 23.
Visual Relation Tasks
Visual relation tasks focus on capturing relationships among visual objects, consist of 12 distinct tasks. We categorize them into three splits: fashion-relation, dual-image relation, and spatial/temporal. The fashion-relation split focuses on fine-grained attribute understanding in clothing images, using the cloth-counting and three-/four-image variants of Fashion200K (han2017fashion200k, ) as well as the query–reference caption split of FashionIQ (wu2021fashioniq, ). The dual-image relation split evaluates comparative reasoning across separate images, drawing on NLVR2 (suhr-etal-2019nlvr, ), CIRR (liu2021cirr, ), the two- and four-image true/false subsets of VISION (bai2023vision, ), and MagicBrush (zhang2023magicbrush, ). The spatial/temporal split targets positional and temporal dependencies, combining single-frame spatial reasoning in VSR (liu2023vsr, ) with temporal understanding tasks from SEED-Bench-2 (li2024seed2, ).
Multi-modal Reasoning Tasks
Multi-modal reasoning tasks require the integration of visual cues with common-sense knowledge, comprising 12 distinct tasks. We include four tasks from IRFL (yosef-etal-2023-irfl, ) that examine the figurative interpretation of images paired with non-literal language. We also include five additional tasks derived from the positive–negative image groups in Bongard-HOI (jiang2022bongardhoi, ) and Bongard-OpenWorld (wu2024bongardopenworld, ). These tasks require fine-grained discrimination between contrasting sets. We incorporate three reasoning challenges from COMICS (iyyer2017comics, ), MIT-States (isola2015mitstate, ), and VizWiz (gurari2018vizwiz, ), all sourced from DEMON (li2024demon, ).
VQA Tasks
The VQA sub-group includes 16 diverse VQA tasks that differ significantly from LLaVA’s pre-training datasets. We include novel question types, such as image-quality queries in Co-Instruct (wu2024coinstruct, ), textbook-style diagram interpretation in TQA (kembhavi2017tqa, ), DVQA (kafle2018dvqa, ), and knowledge-grounded QA tasks from SEED-Bench-2 (li2024seed2, ). We also incorporate unfamiliar image domains, including IconQA (lu2021iconqa, ) and WCVQA (winata2024wcvqa, ). We partition Co-Instruct by question type (e.g., Yes/No, How/what questions), IconQA by difficulty and question category (e.g., Multi-choice or Short answering, kindergarten-level or grade 1-level questions), while we divide WCVQA using their original split.
Client 1 Client 2 Client 3 Client 4 Client 5
$\bullet$ Fashion200K (han2017fashion200k, ): ClothCombination
$\bullet$ FashionIQ (wu2021fashioniq, )
$\bullet$ VISION (bai2023vision, ): 2_Img
$\bullet$ MagicBrush (zhang2023magicbrush, )
$\bullet$ Co-Instruct (wu2024coinstruct, ): 2_ImgCompare
$\bullet$ SEED–Bench–2 (li2024seed2, ): KGQA
$\bullet$ IconQA (lu2021iconqa, ): ShortAnswerEasy
$\bullet$ WCVQA (winata2024wcvqa, ): DishName
$\bullet$ Co-Instruct (wu2024coinstruct, ): HowWhat
$\bullet$ DVQA (kafle2018dvqa, )
$\bullet$ IconQA (lu2021iconqa, ): ShortAnswerHard
$\bullet$ WCVQA (winata2024wcvqa, ): ContextDishName
$\bullet$ Co-Instruct (wu2024coinstruct, ): 3_ImgCompare
$\bullet$ SEED–Bench–2 (li2024seed2, ): InstanceQA
$\bullet$ IconQA (lu2021iconqa, ): MultiChoiceEasy
$\bullet$ WCVQA (winata2024wcvqa, ): AdvContextDishName
$\bullet$ Fashion200K (han2017fashion200k, ): ColorConsistency
$\bullet$ NLVR2 (suhr-etal-2019nlvr, ): Subset1
$\bullet$ VISION (bai2023vision, ): 4_Img
$\bullet$ VSR (liu2023vsr, )
Client 6 Client 7 Client 8 Client 9 Client 10
$\bullet$ Fashion200K (han2017fashion200k, ): StyleConsistency
$\bullet$ NLVR2 (suhr-etal-2019nlvr, ): Subset2
$\bullet$ CIRR (liu2021cirr, )
$\bullet$ SEED–Bench–2 (li2024seed2, ): TemporalQA
$\bullet$ IRFL (yosef-etal-2023-irfl, ): MetaphorSimileMatching
$\bullet$ Bongard–HOI (jiang2022bongardhoi, ): ActionDetection
$\bullet$ VizWiz (gurari2018vizwiz, )
$\bullet$ MIT–States (isola2015mitstate, )
$\bullet$ IRFL (yosef-etal-2023-irfl, ): FigurativeVerification
$\bullet$ IRFL (yosef-etal-2023-irfl, ): IdiomMatching
$\bullet$ Bongard–OpenWorld (wu2024bongardopenworld, ): ConceptDetection
$\bullet$ COMICS (iyyer2017comics, ): Dialogue
$\bullet$ IRFL (yosef-etal-2023-irfl, ): PhraseSelection
$\bullet$ Bongard–OpenWorld (wu2024bongardopenworld, ): ConceptQuery
$\bullet$ Bongard–HOI (jiang2022bongardhoi, ): ConceptQuery
$\bullet$ Bongard–HOI (jiang2022bongardhoi, ): ActionIncoherence
$\bullet$ Co-Instruct (wu2024coinstruct, ): YesNO
$\bullet$ TQA (kembhavi2017tqa, )
$\bullet$ IconQA (lu2021iconqa, ): MultiChoiceHard
$\bullet$ WCVQA (winata2024wcvqa, ): Location
Table 23: Per-client task configuration of DRAKE. DRAKE consists of 10 clients, each with 4 distinct tasks. Client 1, 5, 6 tackle visual relation tasks, Client 7, 8, 9 handle multi-modal reasoning tasks, and Client 2, 3, 4, 10 focus on VQA tasks, as illustrated in Fig. 12.
Unseen Tasks
To assess generalization capability, we add 7 additional tasks that are disjoint from the clients’ tasks in DRAKE. We selected challenging benchmarks requiring extensive training, which are far from fast adaptation. The unseen tasks include novel task types (DreamSim (fu2023dreamsim, ), ImageCoDe (krojer-etal-2022-imagecode, )) and novel visual domains with familiar task format (RecipeQA (yagcioglu-etal-2018-recipeqa, ) subsets proposed in DEMON (li2024demon, ), and HQ-Edit hui2025hqedit ). In addition, we split the Contrast-Caption subset of Mantis (jiang2024mantis, ; yu2022coca, ) into two tasks: one with images overlapping LLaVA’s pre-training distribution and another with previously unseen images.
Figure 12: Detailed configuration of the proposed DRAKE benchmark.
Figure 13: Data samples from DRAKE.
A.19 Limitations and Future Work
While we evaluate FedMosaic across various multi-modal FL benchmarks, e.g., HFLB and our proposed DRAKE, and LLM benchmarks, e.g., Fed-Aya and Fed-LLM, computational constraints limit our experiments to a maximum of 10 clients. Exploring performance with a larger number of clients would better reflect real-world FL scenarios and represents an important direction for future work.
A.20 Impact Statements
This work focuses on federated learning, which enables model training in a distributed manner, eliminating the need to share raw data and thereby mitigating data privacy risks. However, since we assume that the model is trained under distribution shifts, there is a potential for the unintended amplification of issues such as model bias and ethical misalignment. We are committed to taking all necessary measures to address these risks, although tackling these concerns is not the primary focus of this work.
A.21 Detailed Algorithm of FedMosaic
Algorithm  1 provides a pseudocode for the proposed FedMosaic framework. We further provide detailed algorithms for initializing adapters (Algorithm 2) and aligning adapters across heterogeneous clients (Algorithms 3 and 4).
# Input
#     Number of clients $N$, Number of client model types $K$,
#     Set of frozen pre-trained models $\mathcal{W}=\{W_{1},...,W_{K}\}$,
#     Mapping of client to model $V:\{1,...,N\}\rightarrow\mathcal{W}$, Number of rounds $R$,
#     Batch size $B$, Learning rate $\eta$, Gradient frequency $f$, EMA ratio $a$,
#     Small-scale pre-trained model $\mathcal{W}_{S}$, Regularization coefficient $\lambda$
#     Training data streams of $T$ tasks for $N$ clients $\{\mathcal{D}_{1},\mathcal{D}_{2},\dots,\mathcal{D}_{N}\}$,
#     where $\mathcal{D}_{i}=\{\mathcal{D}_{i}^{1},\mathcal{D}_{i}^{2},\dots,\mathcal{D}_{i}^{T}\}$, Number of PQ-LoRA blocks $N_{B}$,
#     Public dataset $\mathcal{D}_{p}$, Conventional-LoRA, PQ-LoRA
E = InitializeAdaptersForModels($\mathcal{W}$, $N_{B}$, $K$) # (Alg. 2)
E = AlignAdapter($\mathcal{W}$, E, $\mathcal{D}_{p}$, $K$, $\lambda$) # (Alg. 3, Alg. 4)
# Initialization for each client
for i in range(N) do
      M[i] = [] # Initialize episodic memory
       # Initialize local and global adapters with the aligned adapter E
       L[i] = E[i]
       G[i] = E[i]
      # Initialize gradient vector with zeros
       g[i] = torch.zeros(d_[len(V(i))])
      # Initialize gating parameters for each layer in L[i]
       beta[i] = [0.0 for _ in range(len(L[i]))]
for task t in range(T) do
       for round r in range(R) do
             for client index i in range(N) do
                   W_i = V(i)
                   g_t = []
                   for j, (x_j, y_j) in $\mathcal{D}_{i}^{t}$ do
                         M[i].append((x_j, y_j))
                         x, y = random_choice(M[i], B)
                        # Compute loss with mixed local-global adapter
                         loss = CrossEntropy(Forward(W_i, L[i], G[i], beta[i], x), y)
[…]
                         if i % f == 0 then
                               loss = CrossEntropy(Forward($\mathcal{W}_{S}$, x), y)
                               g_t.append(LastLayerGrad($\mathcal{W}_{S}$, loss)
                  g[i] = (1 - a) * g[i] + a * Average(g_t)
            # Aggregate global adapters using SIMA
             G = SIMA(L, g)
Algorithm 1 FedMosaic
// Initialize the list of adapter lists for each model
E = []
for k in range(1, $K$) do
       // Initialize adapter list for model $\mathcal{W}_{k}$
       E_k = []
      // Block size per model (PQ-LoRA assigned)
       B_k = len($\mathcal{W}$[k]) // $N_{B}$
      // Remaining layers not evenly divisible
       B_r = len($\mathcal{W}$[k]) % $N_{B}$
       for l in range(1, len($\mathcal{W}$[k])) do
[…]
                   // Append PQ-LoRA
                   E_k.append(PQ-LoRA)
            else
                   // Append conventional LoRA
                   E_k.append(Conventional-LoRA)
      // Add adapter list to full collection
       E.append(E_k)
return E
[…]
# Define hook functions to capture outputs during a forward pass
Function hook_fn(module, input, output):
       self.lora_outputs.append(output)
### Combine PQ-LoRA with the pre-trained model
E_indices = []
for i in range(K) do
       E_index = []
       for j in range(len(W[i])) do
             if E[i][j] == PQ-LoRA then
                   # Attach PQ-LoRA to pre-trained model layers
                   W[i][j] = Attach(E[i][j], W[i][j])
                   E_index.append(j)
      E_indices.append(E_index)
### Align PQ-LoRA A matrices
W_p = W[0] # Set pivot model
W_p.freeze() # Freeze pivot model
for i in range(1, K) do
       W_i = W[i]
       ## Step 1. Attach hooks to LoRA A and perform a forward pass
       # Register hooks to capture the outputs of LoRA A matrix at PQ-LoRA attached layers
       for j in E_indices[0] do
             W_p[j].A.register_forward_hook(hook_fn)
      for j in E_indices[i] do
             W_i[j].A.register_forward_hook(hook_fn)
      ## Step 2. Alignment of A using MSE loss
       total_loss = 0
       for (x, y) in $\mathcal{D}_{p}$ do
             W_p.forward_with_lora_scaled(x, scale=0)
             W_i.forward_with_lora_scaled(x, scale=0)
            total_loss = 0
            # Compute MSE loss for each layer using outputs captured by hooks
             for j in range(len(lora_outputs[i])) do
                   # Compute MSE between LoRA A outputs of W_i and W_p at layer $j$
                   loss_mse = MSE(W_p.lora_outputs[j], W_i.lora_outputs[j])
                   # Compute Regularization loss to prevent deviation from orthogonality (Eq. 23)
                   loss_reg = Reg(W_i[j].A)
                   total_loss += loss_mse + $\lambda$ loss_reg
[…]
      ## Step 3. Post-process A to enforce Orthogonality
       for j in E_indices[i] do
             # Use Singular Value Decomposition to get the closest orthogonal matrix
             U, S, Vt = SVD(W_i[j].A)
             W_i[j].A = U*Vt
Algorithm 3 AlignAdapter(W, E, $\mathcal{D}_{p}$, $K$, $\lambda$) - Part 1
### Align PQ-LoRA B matrices
for i in range(1, K) do
       W_i = W[i]
       ## Step 1. Attach hooks to LoRA B and perform a forward pass
       # Register hooks to capture the outputs of B in PQ-LoRA attached layers
       for j in E_indices[0] do
             W_p[j].B.register_forward_hook(hook_fn)
      for j in E_indices[i] do
             W_i[j].B.register_forward_hook(hook_fn)
      # Forward pass to collect outputs of the LoRA B matrices for CCA
       for (x, y) in $\mathcal{D}_{p}$ do
             W_p.forward_with_lora_scaled(x, scale=0)
             W_i.forward_with_lora_scaled(x, scale=0)
      ## Step 2. Apply CCA to align B matrices
       for j in range(len(W_i)) do
             # Extract B matrix outputs from hooks
             X_p = W_p.lora_outputs[j]
             X_i = W_i.lora_outputs[j]
             # Compute CCA projection vectors
             $\Pi$_p, $\Pi$_i = CCA(X_p, X_i)
             # Initialize B using the CCA transformation
             W_i[j].B = (($\Pi$_i)-1)T · ($\Pi$_p)T · W_p[j].B
E_align = []
for i in range(K) do
       for j in range(len(W[i])) do
             if E[i][j] == PQ-LoRA then
                   E[i][j] = W[i][j]
      E_align.append(E[i])
return E_align
Algorithm 4 AlignAdapter($\mathcal{W}$, $E$, $\mathcal{D}_{p}$, $K$, $\lambda$) - Part 2
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: In the abstract and introduction, we propose two main contributions: (1) SIMA, a novel local model aggregation method designed to reduce interference, and (2) PQ-LoRA, a dimension-invariant adapter that facilitates knowledge sharing across heterogeneous architectures. Additionally, we summarize these contributions clearly at the end of the introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discussed limitations in Sec. A.19.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We provided theoretical proof in Sec. A.1 and Sec. A.2.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide implementation details in Sec. A.5.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: We will release our code after publication.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provide the experimental setups and test details in Sec. A.5.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We ran each experiment with 3 random seeds, and reported averages and standard deviations.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide the details of computational resources we used in Sec. A.5.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We checked the the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discussed impact statement of our work in Sec. A.20.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Although we proposed a new benchmark, i.e., DRAKE, it is composed of a combination of existing publicly deployed datasets that have already undergone safeguards, and no newly scraped datasets are included.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We followed the implementations of methods proposed by their respective authors and cited them accordingly.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We provide the details of our proposed DRAKE and will prepare a documentation and license before releasing the dataset after publication.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: [N/A]
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: [N/A]
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: We use MLLMs as a model, and the training details and the model versions are provided in the appendix.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.